A Bibliography of Publications in Science and Public Policy: 1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.02

Title word cross-reference

1979 [GK79]. 1980s [Ano79e, Bow79].

20-country [BC77].

30 [Mac75].

Aborigines [Ben74]. absence [Ano75-74]. Academic [Per77]. Academy [Ano75-49, Ano75-86]. accelerator [Mot77]. acceptable [Rav79]. accident [Eva76]. accounting [Bro76]. achievements [Hil76, Soe78]. across [Tha76]. act [Ano76-36]. Activities [Ano74-39, Ano74-40, Ano74-41, Ano76-39, Ano74-42, Ano75-77, Ano75-78, Ano75-79, Ano76-37, Ano76-38, Ano77-36, Ano77-37, Ano78x, Ano78y, Ano79j]. Ad [Ano74a, Ano79a, Aur74b, Aur74c, Aur74a]. additives [Col79]. Adhoc [Aur74d]. Adjustment [Ric76]. Administration [Ano76c]. Adult [Ano75a]. Aeronautics [Gab74]. affect [Rob75b]. afford [Ano77d]. Africa [Ano75-50, Ano79k]. African [Leg77]. age [GW75, Leo78]. agent [Ram79b]. agreement [Ano75g]. Agricultural [Abe75, Soe78, oC77]. Agriculture [Ano75b, Ano75c, Ano75o, Ano75-76, Ano77g, Ano78-31, FB79, KA76, Oja75]. ahead [Gau75]. Aid [Ano78a, Ano77-41]. air [Izm74, Mor79]. Allocations [Ano75-82]. alternative [Ano76-35]. America [BC79a]. among [Sag75a]. Analysis [vHF79, Ano77b, Hoe79a]. analyzers [vHF79]. Animal [Ren79, FB79, KD77]. Annual [Ano74-45, Ano79d]. answers [Gol79b]. Application [Ano77a]. Applied [Mor78a, Mac79]. Approach [Azi76, Flo77, Ano75-71, Ber78a, Cor75, FB79]. Appropriate [McR79, BI75, Tei78]. Appropriateness [Ber78c]. Arab [Ami77, Ano77a, Ano77z, ES76, Nas78, Yaa75, Zah74]. arid [Ano76y]. arms [Ano78v]. army [Obu79]. arrangements [Odh79]. artist [Soy78]. arts [Arc78]. aspects [Mab78]. assessing [NP79]. assessment [Ano77-39, Ano78w, Bra77b, Che77, Eis79]. assist [Ano76-34]. assistance [Ric76]. ASTEC [Ano75d]. Attitude [vM77]. Attitudes [KS78, Ano77-29, KB78]. Aurum [Ano74a]. Australia [Ano74-36]. Australian [Ano76c, Ano75d, Ano76c, Ron75, Ron77]. automated [vHF79]. automation [Leo78]. averages [Ano77n].

background [AF75]. Backlash [Ano75e]. balance [Sta77]. Bank [Ano77-41]. barefoot [Ano75j]. basic [Mar76]. Basin [Dic79]. be [Ano77y]. before [Wri78]. Benefit [Ano77b]. Benefit-cost [Ano77b]. Benefits [Ano79b]. Benjamin [Ano76a]. Better [Ano75f, Bla79]. between [Zab78]. Bhargava [Mac75]. Big [Bla79, Tow74]. biological [Ano76-32]. biologist [Gol78a]. biologists [Ham75]. blind [Mac78b]. body [Ano78l]. boffins [Wil79a]. Book [Mac75]. Bracknell [Ogi76]. brains [Dyo77]. Brasil [Ano75-67, Ano76f]. Brazil [Ano74-37, Ano75g, Ano77c]. bread [Ano75-87]. Briefing [Ano74b]. Britain [Ano75-74, Kea77, Law79]. Budget [Ano75-82, Ano74s, Ano76t, Ano77m]. budgeting [Ano79j]. burns [Ano77p]. business [Ano77-39, Fry75, Tow74]. businesses [BL78]. cake [Wil78b]. Call [Ano75h]. can [Ano77d, Ano77y]. Canada [Ano78b, AF75, Dru74]. Canadians [God75]. cancer [Ano78q]. Capital [Ano77d]. care [Jef76, Mar76]. Caribbean [Ano76b]. Case [Boe77, BI75, Col79, Coo77, Sta77]. catastrophe [Bof76]. ceilings [Ano76-32]. census [KA76, Wri78]. cent [Ano74-45]. Centralism [Ano76c]. centres [Nas78]. centuries [Tha76]. Chagas [Pet77a]. Challenge [Cas78, Leg77, Ree77, Ust77]. change [DD74, Ken74, Ram79b, Rob75b, Rob75a, Rot79b]. changes [Szc77]. cheaper [vD79]. chemicals [Ano77-35]. chemistry [vHF79]. children [Rod78]. China [Ano75b, Ano75k, Ano75j, Ano75i, Ded75, SK75]. Chinese [Azi76]. Civil [Cou77]. classics [Ano75-83]. Climate [Rob75a]. Climatic [Mab78, Rob75b]. clinical [vHF79]. Club [Sie75]. Coast [Kan77]. code [Ano75r]. collaboration [Bar75b]. collective [Ust77]. commercial [Coo77]. Commission [Ano76c, vM77, Ano76-34, Ron75]. Commonwealth [Ano76d, Ano76e, God78, Ram79b]. Communicating [Mik79]. Communication [Cle75, Tho74]. Community [Ano75l, Gol79a, Ano75-64]. Comparative [Ano79c]. competition [Ano79f]. competitive [Pri74]. computer [Tit79]. Conception [Law79]. Concern [Tei78]. conference [Arm75]. conflict [NP79]. Consensus [NP79]. consequences [Ano79q]. conservation [Ano74-38, Maz78, Pen78, SK75]. conserver [Ano78b]. Constraints [Lav78, Hoe79a]. context [Tho74]. continuing [Odh79]. Contribution [Sal77a, HE79, Sud79]. Control [Ano76f, Bra77a, Cra76, Lal74, Pet77a]. Converting [Mor79]. cooperation [Ano76e, ES76, Sag75a, Tri76]. corporations [Ada74]. correspondent [Ano75-71, Ano79s, Fis77]. cost [Ano77b, Ano78v]. Costa [Obu79]. costs [Ano76s, KD77]. Council [Ano74c, Mac75, Ano74t, Ano74-43, Ano75d, Ano75-65, Ano75-85, Ano77r, AF75]. Councils [Ano75-29]. Countries [Ded75, Mou74, Abe75, Ada74, Ami77, Ano75-70, Ano77g, Ano77z, Ber78b, Ber78c, Fen76, God78, Krö78, Mor78a, Ric76, Sag75a, Wei79, dH75, vT76]. country [BC77, Obu79]. course [Per77]. creation [Per77]. Crisis [Ano77e, Sal77b, Ano75o, Car76, May75]. Cultural [Mac75, Nse77]. culture [Fre78]. Curiosity [Ano75m]. current [Arm75]. curriculum [Ren79]. cutting [Wil78b]. Cynics [Pen78].

D [Ano75-63, CA74]. dangers [Ano79b]. Danish [Ano74c]. darkness [Ano74n, Car74]. debate [Ice77]. debt [Krö78]. decade [Wil76]. decision [Ano77-35, Lom77]. declaration [Ano75n]. Deficiencies [Kot74]. Defining [Bea76]. Defoliation [Ano74d]. Defunct [Kin75]. degradation [Dub76]. Delhi [Mac75]. Delusions [Mot75]. demand [Hae78]. Demands [Nic74]. Democratic [Ano79r, Ano76h]. Democratization [Sab79]. dependent [Ano74-29]. desertification [Mab78]. deserts [Ano78z]. design [Ber78b]. developed [Ric76]. Developing [Ded75, Mou74, Abe75, Ada74, Ano75-70, Ano77g, Ber78b, Ber78c, BL78, Fen76, God78, Krö78, Mor78a, Wei79, vT76]. Development [Ano75-85, Azi76, DD74, GK79, Pro76, AN75, Ano75-67, Ano76y, Ano78z, BI75, Che77, Hoe79a, HE79, Kea77, Kot74, Lal75, Nse77, Rah79, Ram79a, Ram77a, Sud79, Ver77, WL79, Wil76]. died [Ano75-88]. different [KS78]. diffussion [Loc78]. Digest [Ano77-28, Ano77-27]. dilemma [Ano76b, Bor74]. diminishing [Gia79]. Disciplines [Ano76g, Ano77f]. disease [Pet77a]. disjunctive [Bar77]. dissemination [Gol78b]. do [Ano75-89, GW75, Mey78]. doctor [Ano75j]. doctrine [Ded78]. Doubts [Mor78b]. Draft [Ano75n]. dreams [Rot79a]. drought [Ano75-66]. dukes [Lar77]. Dutch [Ano79d, Boe77]. dynamic [Ano75-46].

East [Ano78c]. ecological [Mab78].

economic

[Ano75-72, Bog77, Boj79, Lom77, Vai77]. Economics [Gia79]. economy [Ano75-50, Ano78-32, Kin78, Pri74]. ecosystem [Ano79q]. ecosystems [Ano79t]. Editorial [Ano74e, Ano74f]. Education [Ano74g, Pet77a, Ano75a, Pro76]. Educational [Ano77f]. EEC [Ano75p]. effect [Rob75a]. effectiveness [Kea77]. Effects [Maz78, Hod77]. Einstein [Cla78, Whe79]. electricity [Mey78]. elite [Kin74]. emigré [Tor79]. Employee [Ano76h]. Employment [Leo78, Muh76, Abe75]. Encounter [Ano74-43]. Energy [Ano75o, Ano75p, Ano76i, Ano77g, Ano78d, Bro76, Mey78, Mou74, Nic74, vD79, Ano74-29, Ano74-38, Ano75h, Ano76-33, Ano78o, Ano78-31, Fel75, Hae78, KS78, Maz78, Mum74, Orr77, Pen78, Tel77]. energy-dependent [Ano74-29]. engineering [Ano75-30, Ano79b, Ber78b, Law79]. engineers [Alb77]. England [Wri78]. Enough [Mum74, Ano76q]. enterprise [Dun74, Hor79]. entrenchment [Col79]. Environment [Ano75q, Ano77-35, Dar75, Eld79, Fai77, Sem77, Wei77]. Environmental [Ano79e]. Epistemology [Arc78, Mar74]. era [Ami77]. Erosion [Cra76]. establishment [Izm74]. Estimating [Wri78]. Ethical [Ano75r]. ethics [Gel77]. Europe [Ano75l, Ano76j, Ano77-38, Ano78c, Ano78e, BC79a, Car79, Col77, Gol77b, Mot77, Tri76].

European [Ano75-64, Ano76-36, Ano78-31, Ano79f, ES76, Flo77, Joh77, vM77]. evaluation [Eva76]. evolution [Cro78, Gol78a, Tit79]. expectation [Eis79]. experience [Bro78, Cle79, Eis79]. Experiment [Ver77]. experts [Ano76d]. Exploring [Rav78].

fact [Fel75]. Facts [Ano74h, Ano74i, Ano74j, Ano74l, Ano75s, Ano75t, Ano75u, Ano75v, Ano75w, Ano75x, Ano75y, Ano75z, Ano75-27, Ano76k, Ano76l, Ano76m, Ano76n, Ano77h, Ano77i, Ano78f, Ano78g, Ano79g, Ano74k, Ano79h]. farms [Rod78]. fate [Tal79]. fears [Mor78b]. Federal [Ano78-30, Ano75a, Fry77a]. feed [Szc77]. field [BC79b]. fields [ES76]. Figures [Ano74h, Ano74i, Ano74j, Ano75s, Ano75w, Ano75x, Ano75y, Ano75z, Ano75-27, Ano76l, Ano76m, Ano76n, Ano78g, Ano74k, Ano74l, Ano75t, Ano75u, Ano75v, Ano76k, Ano77h, Ano77i, Ano78f, Ano79g, Ano79h]. financial [Nas78, Ram79a]. Financing [Bir76, Fry77b, Jon77]. first [Wri78]. Fish [Ano75-28, Cra76]. flared [Ano75-48]. flight [Ano75-89]. floods [Ano75-66]. flow [Boj79]. Flying [Mac78b]. Focus [Ded75]. Food [Ano77j, Ano75-92, Bea76, Bor74, Car76, Rob75b, Rob75a]. force [Ano75i, Ano76-34, BC79b]. forecast [Ano76o]. forest [Ano75-47, Ano79t]. Forestry [Dar75]. forests [Tal79]. forms [KS78]. Forthcoming [Ano74m]. forward [Mor79]. Foundation [Col75, Flo77]. Four [MPG74]. fragments [Fre78]. Framework [Haw77]. Franklin [Ano76a]. freedom [Per77]. French [KB78, oB77]. fresh [Cra76]. full [Lar77]. fund [Ano76e]. Funding [Ano76o, Dru74]. Future [Ano74g, Ano74n, Ano78-33, Car74, Ano77d, Ano78p, Ano78w, Dun74, Gar75, Ham75, HLRS76, Muh76]. Futures [Rob78].

G.P. [Jef76]. gain [Ano75-89]. gap

[Yaa75, Zah74]. gas [Ano75-48]. General [dC78a, Ano75-29]. generated [vD79]. Generating [Ano78h]. genetic [Ano79b]. German [Ano79r]. Germany [Ano75a, Ano75g, Law79]. get [Wil78b]. Getting [McQ77b]. giant [Mot77]. girl [Ano76z]. given [Ano77y]. Global [Las75, Ano76s, Tho74]. God [Cla78]. going [Ano77l]. Government [Ano75-30, Ano76c, Hag79a, Lom77, Ano77-39, Gol78b, Hod77, Rob78, Fry77a]. Grain [vT76]. Grandeur [Mot75]. Great [Ano75-74]. Green [Rot79a]. group [Ano76d]. Growth [Ano74-45, Ano75-46, Ano78-30, Ave75, For76, Nas78, Ogi76]. guaranteed [Ano77p]. Guidelines [Sag77a]. Guinean [BB77].

Hague [Ano79s]. Hazard [Eis79]. hazards [Ano77l, Sem77]. Health [AN75, Ano77k, Coo76, Hor79, Mah76, Mar76, Sem77, Ano75-50, Ano77l, Ano77-41, Ano78a, Ano78r, Bra77a, God75, Lin76, Pet77b]. heroism [dSP78]. Highlights [Ano79d]. Hoc [Aur74c, Ano74a, Ano79a, Aur74a, Aur74b]. Homo [Ram79c]. hot [Mor79]. hour [Ano77p]. House [Ano76a]. Human [Ram77a, Ano74r, Ano77k]. Humanities [Ano77m]. humanity [Oja75].

idea [Mad78]. II [dC78a]. imperialism [Bud75]. Implications [Pan77, Ano74o, BB77, Vai77]. importance [Oja75]. Impressive [Ano77n]. Increase [Ano76o]. Increasing [Kea77]. independence [BB77]. Index [Ano74p, Ano75-31, Ano76p, Ano77o, Ano78i, Ano79i]. India [Ano75-66, Ded75, Mac75]. Indian [Mac75, Ano77e]. indicators [Ano78j]. individual [Ust77]. Industrial [Col77, Gol77a, Kat76, Ano76-41, Fry77a, Lav78]. industrialized [Cro78]. Industry [Ded75, Fry77a, Ano74-38, Hag79a, Mad75, Pan77]. inflation [Ano77r]. Information [Jan78, Ano78q, Boe77, Hag79b, Mik79]. Informational [DRR78]. innnovation [Loc78]. innovation [Col77, Fry77a, Gol77a, Haw77, Lav78, vHF79]. innovative [Fry77b, Jon77]. insights [Tor79]. Institute [Dyo77, Ano75-86]. Institution [Kin75]. institutional [Odh79]. institutions [Ano77f, Nas78, Ram79a, Wei76]. instruments [Sag79]. integrating [Tor78]. intellectual [Gol79a]. Intelligence [Mac79, Ded78]. interdependence [Ano74-28]. Interdisciplinary [Win74, dC78b]. interest [Bra77a]. International [Ano74q, Ano78j, Ano78k, Tri76, Ano75-72, Bar75b, Bog77, Boj79, DRR78, Tin76, dC78b, GK79]. intervention [Hod77]. Introduction [GK79]. investment [Ano74r, Ano78d]. investments [Abu77]. involvement [Ano77-30, Dic79, Nas78]. Irish [Ano78s]. Isolation [RS77]. Israeli [Kat78, Lav78, Soe78, Tor78, Yaa75, Zah74]. ISSTA [GK79]. issues [Bof76, Ram77b]. Ivory [Kan77].

January [GK79]. Japan [Ano74-29, Fai77]. Jordan [Ano78l]. journal [Gar75]. journalism [Ver77].

key [Ano74-29]. kids [Wil79a]. Kingdom [Cou77]. Knowledge [AF75, Gol78b].

land [oC77]. Language [Pin78, Wil77]. larger [Whe79]. last [Gra78]. law [Bar75a]. lay [Ano77p]. LDCs [Ded78]. lead [Ano75-90, Col79]. leap [Bar77, Tha76]. Learning [Bro78]. Lebanon [Hoe79b]. less [Ano75f]. Let’s [Wil78b]. level [Ano77q]. life [Abr75, Ahm77, Ano76-31, Gab74, Mum74]. Limits [For76, Ave75]. link [Ndo78]. lived [Ano75-88]. livestock [Szc77]. livestock/feed [Szc77]. Lomé [Ano75-90]. Long [Ano76s]. Long-term [Ano76s]. look [Ano74s, Ano74t, Lal75]. lost [oB79]. Low [Ano77q]. Low-level [Ano77q].

M [Mac75]. machine [Zab78]. make [Ano76r]. making [Ano77z, Ano78l, Lom77]. Malaria [Ano74r]. man [Dar75, Ham75, Mum75, Tit79, Zab78]. management [Ano78z]. mankind [Mac78a]. Manpower [HLRS76, Ano75-84, Une79]. many [Alb77]. martyrs [Pen78]. Materials [Ano75-64, CA74]. matter [Ano75-32]. Measuring [Jef76]. meat [Ano77s]. mechanization [Abe75]. Medical [Ano75-85, Ano77r, ATM77, Ano75-84, Gel77]. Meetings [Ano74m]. mercury [GW75]. Metamorphosis [Pri77]. methane [Ano78h]. Methodology [Fry79]. Mid [Wil76]. Mid-term [Wil76]. Migration [God78]. mile [Ano77p]. Milk [Fen76]. minerals [Mic78]. minister [Ano75q, Ano76t]. Ministry [Kan77]. missing [Ndo78]. missionaries [ATM77]. Mobilizing [Wei79]. model [Ano75-46]. modified [Ano75-46]. money [Ano77p, Coo77]. Mongolia [Tha76]. motion [Mor79]. much [Mey78]. Multinational [Ada74, Dun74]. Multinationals [Lal75]. Multistat [Tit79]. myth [Fel75]. myths [Hil76].

N [Mac75]. National [Ano75-65, Ano79j, Ram79a, Ano75-71, Ano75-67, Ano78l, Nse77]. NATO [Sud79]. Near [Ano78-33]. necessity [Ano76h]. need [Mey78, Tel77]. needed [Ano76-42]. needs [Ano77d, Dar75, dSP78]. Neglected [Ano75-33]. Netherland [Ano76t]. Netherlands [Ano74s, Ano75-71, Ano76i, Ano77m, Ano78-28, Ano78-27, Cor75]. networks [DRR78]. new-style [Ano76t]. News [Ano74q, Ano74u, Ano74v, Ano74w, Ano74x, Ano74y, Ano74z, Ano74-27, Ano75-34, Ano75-35, Ano75-36, Ano75-37, Ano75-38, Ano75-39, Ano75-40, Ano75-41, Ano75-42, Ano75-43, Ano75-44, Ano75-45, Ano76u, Ano76v, Ano76w, Ano76x, Ano77t, Ano77u, Ano77v, Ano77w, Ano77x, Ano78m]. Next [Mum75, Ano74-28, Tit79]. NGOs [Mor79]. Nigerian [Cle79]. no [Mah75]. noise [RG77]. non [Ano74o]. non-profit [Ano74o]. Nordic [dH75]. Normal [Ano75-46]. North [Ram77b]. North/ [Ram77b]. Norway [Ano79p]. Norwegian [Ano78n]. Notes [Sag75b, Sag75a]. Nothing [Ano77y]. Nuclear [Ano74-29, Ano78o, Ano78p, Eva76, Fel75, Hay77, Hil76, Nic74, Wil79b, Ano75g, Ano76s, Ano77c, Ano77l, Bof76, Bro76, HLRS76, Ice77, Tel77, oB77]. nurturing [Ano75-47]. nutrition [Ano77j, Bea76, Cép76, Pin78].

Observations [Ano75k]. obsolete [Kin74]. Obstacles [Ano76y]. Occupational [Ano78q, Ano78r]. odds [Bof76]. OECD [Ano75-63]. OEeD [Ano78-34]. offshore [Ano75-30]. oil [AS78, Ami77]. Oilfield [Ano75-48]. One [Gol79a]. OPEC [Abu77]. opinion [Ice77]. option [Orr77]. options [MPG74]. order [Ano75-72, Boj79, Rah79, Tin76]. organic [Ano78h]. Organizational [Ano75-49]. oriented [Ano75m]. other [Pan77, Whe79]. our [Ano75-71, Ano79s, Fis77, Rod78]. Outlines [Bog77]. output [Ano76-41, BC77]. ozone [Ano76-43].

Palestinian [DD75]. paper [Ano76r, Kin78]. Papua [BB77]. participation [Ano76h]. past [Ano75-76, Muh76]. Patent [Bar75a]. Patents [Arm75]. Paulo [Ano75e]. Pavane [Kin75]. People [oC77, Ano75-32, McQ77a]. perceptions [Rav79]. performance [Ano78s, Ano79c]. perspective [God75, Wil79b]. Perspectives [Ice77]. petrol [Col79]. phenomena [Rav78]. Physical [Mic78]. Physicians [Ham75]. physics [Pre77]. place [Ano75-50]. plan [Ano75-29, Ano75-67, Ano75-68]. planners [Fry79]. Planning [Ano75-51, Ano78-28, dC78b, Ano79p, Coo76, Pin78, Ron75, Win74, Ano74c]. play [Arm75]. pluralism [Ano76c]. point [Mah75]. points [Rav77]. Poland [Kub78]. policies [Ano75-70, Ano79e, Ano79r, Fai77, Kat76, Leo78, May75, Sag77a, vT76]. Policy [Ano74c, Ano74-37, Ano77z, Ano78r, AF75, Cor75, Dob74, Abr74, Ano74s, Ano74-36, Ano75h, Ano75-64, Ano75-71, Ano75-85, Ano76c, Ano78l, Ano78-28, Ano78n, Ano78-27, Ano79d, Ano79f, Bea76, Che78, Cle79, CA74, Gol77b, Hoe79a, Joh77, Kot74, Loc78, McQ77a, Muh76, Nse77, Sag79, She77, Tor78, Win74, Ron75, dH75]. policy-making [Ano78l]. Polish [Ano75r, Ano75-49]. political [Kin78]. Politics [RG77, Ano76-40]. polities [NP79]. Pollution [Tel77, Ano76f]. Poor [Ano78s]. Popularization [Gol74, GD77, KB78]. Population [Ano74-45, Bor74, McN78, Wri78]. post [Ami77]. post-oil [Ami77]. Poverty [McN78, Tel77]. Power [AF75, Nic74, Ano74-29, Ano77c, Bro76, Eva76, HLRS76, Hil76, Ice77]. pp [Mac75]. practical [Arc78]. Pregnancy [Ano76z]. presents [Ano74s, Ano76t]. President [Col75]. Press [Ano77-27, Ano77-28, Ram79b]. prevention [Ano78q]. price [Bor75, Szc77]. Principles [Izm74]. priorities [Ano75-33, Ano75-75, Coo76]. problem [Krö78, Wil79b]. problems [Ano75-49, Hil76, Ric76]. production [KS78, KD77, Ren79]. professionals [God78]. profit [Ano74o]. program [Dic79]. programme [Ano77c, Odh79]. progress [Ano75c, Ano75-80, Mad78, Sag75b, Sie75]. Progressive [Dub76]. project [Ber78b, Sag75b, Sag79]. projects [Ano77-41]. prospect [Ano77-40]. Prospects [AS78, Ano79e, Ano79k]. prosperity [Ano74-28]. protecting [Loc77]. protein [KD77]. provides [Cra76]. Pruthi [Mac75]. Psychosocial [BB77]. Public [Ano77-29, Ano77-30, AF75, Boe77, Rav79, Ano75-50, Ano78q, Dic79, Ice77]. Publication [BC77]. Publications [Ano74-30, Ano74-31, Ano74-32, Ano74-34, Ano75-52, Ano75-53, Ano75-54, Ano75-55, Ano75-56, Ano75-57, Ano75-58, Ano75-59, Ano75-60, Ano75-61, Ano75-62, Ano76-27, Ano76-28, Ano76-29, Ano76-30, Ano77-31, Ano77-32, Ano77-33, Ano77-34, Ano78t, Ano78u, Ano79l, Ano79m, Ano79n, Ano79o, Ano74-33]. pulp [Kin78]. Purposes [Soe78].

quality [Abr75, Ano76-31, Gab74, Izm74, Jef76, Kat78]. Quantity [Kat78]. Qu’Appelle [Dic79]. quest [BI75]. questions [Ano78p]. Quo [Ram79c]. Qureshi [Mac75].

R [Mac75, Ano75-63, CA74]. R&D [Ano76o, Ano78n, Ano78-30, Hag79a, Kat76]. race [Ano78v]. radioactive [Ano77q]. Rahman [Mac75]. Raising [Ano76-32]. Rasmussen [Bof76]. Rate [Ano74-45]. Rational [Ano76-33]. ratios [Szc77]. Raw [Ano75-64]. RAWB [Ano79d, Ano79s]. read [Bla79]. Real [Ano78v]. really [Mey78]. Recommendations [Odh79, Ano74-44]. refinement [Gra78]. Reflections [Gol78b, Loc77, Sag77b, dC78a]. reform [Bar75a]. regarding [KS78]. regional [Nas78]. regulatory [Ano77-35]. related [Ano79p]. Relations [Mac75, Bog77, Rav77]. Relationship [Zab78]. reliance [Bar75b, Red77, Sag75a]. Report [Col75, Ano79d, Ron75, Sie75, Ano76c]. Reports [Ano74-35, Ano74-43]. Republic [Ano75a, Ano79r]. requirements [HLRS76]. Research [Ano75-65, Ano75-85, Ano77r, Ano78w, Ano79p, Ndo78, Ano75m, Ano75l, Ano75-29, Ano75-33, Ano75-73, Ano75-81, Ano76i, Ano76-42, Ano77m, Cle79, DRR78, Mor78a, Rob78, Soe78, Wei76, dC78b, Kan77]. resolution [NP79]. resolutions [Ano75q]. Resource [KD77, Gra78]. resources [CA74, Coo76, Une79]. Responsibility [Las75, Lal74]. responsible [Ano75-91]. return [Mah75]. returns [Gia79]. Review [Mac75]. revised [Bof76]. revisited [For76]. Revolution [dSP78]. revolutionary [Ano75i]. revolutions [Sag77b]. Rica [Obu79]. ride [Mac78a]. right [Ano77k]. Rights [Ben74, Mar76, Ram77a, Wil78a]. risk [Fry77b, Jon77, NP79]. risks [Gol77a, Orr77, Rav79]. road [Gau75]. role [ATM77, Ano75-30, Ano75-65, Bra77b, Fry77a, Gol78b]. Romania [Ano75-73]. Rome [Sie75]. room [Lar77]. Royal [Ano76c, Ano76-34]. Rs [Mac75]. runs [GW75]. Rural [Azi76].

S [Ano79r]. saddle [Mac78a]. Safety [Sta77]. São [Ano75e]. saved [vD79]. Saving [Ano75-66]. scarcity [Mic78]. schemes [Fen76]. School [Ahm77]. Science [Ano74-36, Ano74-37, Ano74-43, Ano75d, Ano75-82, Ano75-70, Ano75-71, Ano75-67, Ano75-72, Ano75-69, Ano75-68, Ano76-34, Ano77-35, Ano78-27, AF75, Bir76, Cle74, Cle75, Cor75, Ded75, Dob74, Flo77, Fre78, Gol77b, GK79, Hoe79a, Ken74, Las75, Loc78, Mad75, Mar74, May75, McQ77a, Rah79, Wei77, Ano74s, Ano75k, Ano75-75, Ano76c, Ano76t, Ano77a, Ano77e, Ano77m, Ano77-29, Ano77-30, Ano78j, Ano78l, Ano78p, Ano79d, Ano79j, Bar75b, Boe77, Boj79, BC79b, Bre76, Cou77, Dru74, Fis77, Fry75, Gol74, GD77, Hoe79b, KB78, Kat78, Kub78, Mac79, Mal79, Pri77, Rav77, Sab79, Sag79, Sal77b, She77, Sud79, Tow74, WL79, Yaa75, Zah74, Ron75, dH75, Mac75]. sciences [Ano76g, BC77, Sal77a, Tri76, Ano75-49, Ano75-86]. Scientific [Ano74c, Ano75-73, Bud75, Hag79b, Kan77, Tor79, Une79, Ano75r, Ano75-85, Ano77f, Che78, Gar75, Hor79, Mik79, Mor78a, Sag77b]. scientific-technological [Sag77b]. scientists [Alb77, Tor78, Tor79]. Search [Gol79b]. second [Wil76]. Secretary [dC78a]. Secretary-General [dC78a]. sector [Ano78s]. sectors [Pan77]. security [Ano76-36]. seekers [Whe79]. selection [Ree77]. Self [Bar75b, Gra78, Sag75a, Red77]. Self- [Bar75b]. Self-refinement [Gra78]. Self-reliance [Sag75a, Red77]. semi [Ano76y]. semi-arid [Ano76y]. Seminar [Ano76-35]. Sentiens [Ram79c]. service [AN75, Ano75l, Pet77b]. services [Ano78s]. Sex [Pre77]. shaping [Ano78p]. shortage [Ano75-51]. Sickness [Ano75-74]. Significance [Ano74-38, Ano76-36]. Singapore [GK79, Bek79]. situation [Ano75p]. sixth [Ron75]. skills [Ber78b]. small [BL78, Rot79b]. Smallpox [Mah75]. Social [Abr74, Ave75, Che77, Cle74, Dru74, Lal74, Orr77, Ahm77, Ano75-80, Ano76g, Ano79p, Joh77, Mic78, Rah79, Rav78, Sal77a, Tri76]. Socialism [Bre76]. societies [Cro78]. Society [Las75, Ano74o, Ano78b, Ano78-29, Jan78, Loc77, Mey78, Rav77, Sal77b, Wei76]. Socio [Vai77]. Socio-economic [Vai77]. Sociocultural [Ano79q]. Sociology [Ano75-75, Cép76, Che78]. Solar [Mou74]. South [Ano74d, Ram77b]. Soviet [Ano75-76, Ano75-86, Ano78-29, Dob74, Pet77b, Tor78, Tor79]. space [Ano75-89]. special [Ano79s]. SPF [Ano74-39, Ano74-40, Ano74-41, Ano74-42, Ano75-77, Ano75-78, Ano75-79, Ano76-37, Ano76-38, Ano76-39, Ano77-36, Ano77-37, Ano78x, Ano78y]. stage [Ano74-28]. standards [Izm74]. standing [Law79]. state [Arm75, DD75, Ron77]. Statement [Bek79, Ano76-43]. statistics [Lin76]. steady [Ron77]. step [McQ77b, Tit79]. stock [vT76]. STPI [Sag75b]. STR [Ano75-80]. Strategies [Ano77-38, BC79a, Car79, Gol77a]. Strategy [Ano78z, Bow79, oB77]. Stratification [Cle74]. Structural [Cro78]. Student [Ano75-81]. Studies [Ano74g]. study [AF75, Sta77, Dic79]. style [Ano76t]. subordination [RS77]. substitutability [GW75]. Sudarshan [Mac75]. Survey [Ano79r, BC77]. surveys [Abr74]. sustainable [Jan78, Mey78]. Sweden [Ano74g, Ano75-82, BI75]. Swedish [Ano75-29, Bro78]. Swiss [Ano74-43]. Symposium [GK79, Bek79]. system [Bog77]. Systems [FB79, Jan78].

T [Ano79r]. taboos [Ano77j]. tackles [Ano78e]. taken [Ano77y]. Tank [Ano74-44]. task [Ano76-34]. teacher [Pre77]. Teachers [Ano76-40]. team [DRR78]. technical [Ano76e, Hor79, Mik79, Une79]. technician [HE79]. Technological [Ber78c, BL78, Rot79b, Wil78a, Che78, Hag79b, Pri74, Ram79a, Sag75a, Sag77b, Tri76]. Technology [Ano75-83, Ano77-39, Ber78a, Bow79, Bra77b, Ded75, GK79, Red77, Sea75, Soy78, Wil77, Ano75-72, Ano76-35, Ano77a, Ano77-30, Ano78j, Ano78l, Ano78w, Ano79c, Ano79j, Bar75b, Boe77, Boj79, Bra77a, Che77, Col79, Gia79, Hoe79a, Hoe79b, Mac75, Mad75, Mal79, McQ77b, McR79, Rah75, Rah79, Sag77a, Sag79, Sud79, Tei78, Tho74, WL79, Wei79, Yaa75, Zah74]. techology [Ano75-67]. telephone [Cas78]. Ten [Ano79s, Ano75-76]. term [Ano76s, Wil76]. terrorism [Hay77]. theory [Abr74]. there [Ano76q, Gar75]. Things [Mac78a]. Think [Ano74-44]. Third [AN75, Sag75a, Sea75]. threat [Rah75]. threatened [Ano75c]. time [Ano76r, Ken74]. today [Joh77, Wei77]. Tokyo [Ano75n]. tomorrow [Ano75-28, Joh77, MPG74]. too [Alb77].

Tourism [Eld79, Ano78k]. town [Ogi76]. towns [Ano75f]. tractor [Bor75]. Trade [Ano78-29]. traffic [Ano75f]. training [Alb77, Ber78a]. Transfer [Ber78b, Hoe79b, Mal79, Ano75-83, Ber78c, Ber78a, BL78, Bow79]. transformation [Ano79q, Mum75]. Trends [Ano75-84, Ano75-63, Lin76]. Tropical [Ano79t, Tal79]. tropics [Ano76r]. Turning [Rav77].

UK [Ano74-38, Ano75-85, Ano78d, Mad75]. Ultimate [Ree77]. UNCSTD [Mor79]. underlying [Izm74]. unemployed [Ano77p]. unemployment [Ano78e, Ano78-34]. Unexpected [Hod77]. Union [Dob74, Tor79]. unions [Ano78-29]. United [Cou77]. Universities [Ano74t, Bra77b, Leg77, Ron77, Ron75]. university [Wei76]. Unless [Ano78-33]. unmarried [Ano76z]. Uranium [Ano77-40, Ano75-51, Ano76q]. USA [Ano75-86, Tor79]. Use [Ano78-31, Ano76-33]. USSR [Ano75-81, Ano76-41, Ano78o, Ano78-32, DD74, Izm74, Pro76]. Utilisation [Mou74].

Vadis [Ram79c]. Value [Kub78]. VCM [Sta77]. versus [Mic78]. veterinary [Ren79]. viable [DD75]. Vietnam [Ano74d]. view [Whe79]. viewpoint [Rit74]. views [Fry77a]. VIIth [Ano75-68]. Vision [Ano78-33]. Vojvodina [Ano75-69]. volume [Ano74p]. vs [Ano76c, Fel75]. vulnerability [Gia79].

Wanted [She77]. War [Ano75-87]. waste [Ano77q, Ano77s]. wastes [Ano78h, Wil79b]. Water [SK75]. way [Ano75-88]. weapons [Ano76s]. West [Abu77, Ano75g]. Western [Maz78, Sea75]. Where [Ano76-42, Rod78]. Whizz [Wil79a]. Who [Ano75-91, Gel77]. will [Ano75-90]. Witchcraft [Mar74]. without [Obu79]. WMO [Ano76-43].

Woman [Ano75i]. Women [WL79]. work [Law79, Sem77, Tor79]. workers [Ano75r]. Working [Ano75j, Ano76-31]. World [Ano74-45, Ano75-92, Ano77-41, Car76, Ano75-46, Ano77a, Gol79a, KA76, May75, Pri74, Rob75b, Rob75a, Szc77, oB79, AN75, Sag75a, Sea75]. worldwide [Ice77].

Year [Ano74-43, Ree77, Ust77, Mah76]. years [Ano75-76, Ano79s, Ogi76]. Youth [Ano78-34, Ano78e].

zones [Ano76y].

References

Abercrombie:1975:AME

[Abe75] K. C. Abercrombie. Agricultural mechanization and employment in developing countries. Science and Public Policy, 2(10):442– 447, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Abrams:1974:SST

[Abr74] Mark Abrams. Social surveys, theory and policy. Science and Public Policy, 1(9):250–254, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Abrams:1975:QL

[Abr75] Dr Mark Abrams. The quality of life. Science and Public Policy, 2 (5):212–217, May 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aburdene:1977:OIW

[Abu77] Odeh Aburdene. OPEC investments in the west. Science and Public Policy, 4(2):206–208, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Adam:1974:MCD

[Ada74] Györy Adam. Multinational corporations in the developing countries. Science and Public Policy, 1 (6):125–129, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aucoin:1975:KPP

[AF75] Peter Aucoin and Richard French.

Knowledge, power and public policy: Canada science Council background study 31. Science and Public Policy, 2(2):79– 84, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ahmed:1977:SSL

[Ahm77] Manzoor Ahmed. School of social life. Science and Public Policy, 4(6):517–522, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Alberty:1977:WTT

[Alb77] Robert A. Alberty. Are we training too many scientists and engineers? Science and Public Policy, 4(1):16–21, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Amin:1977:ACP

[Ami77] Mahmud S. Amin. Arab countries in the post-oil era. Science and Public Policy, 4(3):257– 259, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ammundsen:1975:HSD

[AN75] Esther Ammundsen and K. W. Newell. Health service development in the Third World. Science and Public Policy, 2(8):361– 362, August 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:AHA

[Ano74a] Anonymous. Ad hoc by aurum. Science and Public Policy, 1(8): 186–188, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:B

[Ano74b] Anonymous. Briefing. Science and Public Policy, 1(1):9– 12, January 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:DCS

[Ano74c] Anonymous. Danish Council for Scientific Policy and Planning. Science and Public Policy, 1(12): 401, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:DSV

[Ano74d] Anonymous. Defoliation in south Vietnam. Science and Public Policy, 1(6):129, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Ea

[Ano74e] Anonymous. Editorial. Science and Public Policy, 1(1):2A, January 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Eb

[Ano74f] Anonymous. Editorial. Science and Public Policy, 1(2):22, February 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:EFS

[Ano74g] Anonymous. Education for future studies in Sweden. Science and Public Policy, 1(6):118– 121, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:FFa

[Ano74h] Anonymous. Facts in figures. Science and Public Policy, 1(6):114– 115, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:FFb

[Ano74i] Anonymous. Facts in figures. Science and Public Policy, 1(8):210– 212, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:FFc

[Ano74j] Anonymous. Facts in figures. Science and Public Policy, 1(9):255, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:FFd

[Ano74k] Anonymous. facts in figures. Science and Public Policy, 1(10):312– 315, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:FFe

[Ano74l] Anonymous. Facts in figures. Science and Public Policy, 1(11):361– 363, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:FM

[Ano74m] Anonymous. Forthcoming meetings in 1974. Science and Public Policy, 1(9):264, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:FD

[Ano74n] Anonymous. Future darkness 1. Science and Public Policy, 1(6): 98, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:INP

[Ano74o] Anonymous. The implications of a non-profit society. Science and Public Policy, 1(9):227–229, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:IV

[Ano74p] Anonymous. Index to volume 1. Science and Public Policy, 1(1):1– 6, January 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:IN

[Ano74q] Anonymous. International news.

Science and Public Policy, 1(9): 217–226, September 1974. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:MHI

[Ano74r] Anonymous. Malaria and human investment. Science and Public Policy, 1(8):170, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:NBP

[Ano74s] Anonymous. Netherlands budget presents new look for science policy. Science and Public Policy, 1(2):29–32, February 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:NLU

[Ano74t] Anonymous. New look universities Council. Science and Public Policy, 1(10):??, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Na

[Ano74u] Anonymous. News. Science and Public Policy, 1(1):??, January 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Nb

[Ano74v] Anonymous. News. Science and Public Policy, 1(2):??, February 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Nc

[Ano74w] Anonymous. News. Science and Public Policy, 1(6):??, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Nd

[Ano74x] Anonymous. News. Science and Public Policy, 1(8):??, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Ne

[Ano74y] Anonymous. News. Science and Public Policy, 1(10):269–278, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Nf

[Ano74z] Anonymous. News. Science and Public Policy, 1(11):319–324, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Ng

[Ano74-27] Anonymous. News. Science and Public Policy, 1(12):376–386, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:NSP

[Ano74-28] Anonymous. The next stage — prosperity through interdependence. Science and Public Policy, 1(12):366, December 1974.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:NPK

[Ano74-29] Anonymous. Nuclear power:

the key in an energy-dependent Japan. Science and Public Policy, 1(8):188–190, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Pa

[Ano74-30] Anonymous. Publications. Science and Public Policy, 1(2):35– 38, February 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Pb

[Ano74-31] Anonymous. Publications. Science and Public Policy, 1(8):182, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Pc

[Ano74-32] Anonymous. Publications. Science and Public Policy, 1(9):259– 263, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Pd

[Ano74-33] Anonymous. publications. Science and Public Policy, 1(10):311– 312, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:Pe

[Ano74-34] Anonymous. Publications. Science and Public Policy, 1(11):360, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:R

[Ano74-35] Anonymous. Reports. Science and Public Policy, 1(1):16– 19, January 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SPA

[Ano74-36] Anonymous. Science policy in Australia. Science and Public Policy, 1(12):398–401, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SPB

[Ano74-37] Anonymous. Science policy in Brazil. Science and Public Policy, 1(6):117–118, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SEC

[Ano74-38] Anonymous. Significance of energy conservation for UK industry. Science and Public Policy, 1(8): 193–201, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SAa

[Ano74-39] Anonymous. SPF activities. Science and Public Policy, 1(1):15, January 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SAb

[Ano74-40] Anonymous. SPF activities. Science and Public Policy, 1(6): 130, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SAc

[Ano74-41] Anonymous. SPF activities. Science and Public Policy, 1(8):213, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SAd

[Ano74-42] Anonymous. SPF activities. Science and Public Policy, 1(10): 316, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:SSC

[Ano74-43] Anonymous. Swiss Science Council reports on “Year of Encounter, 1973”. Science and Public Policy, 1(6):100–102, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:TTR

[Ano74-44] Anonymous. Think tank recommendations. Science and Public Policy, 1(8):190–192, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1974:WPP

[Ano74-45] Anonymous. World population at 2 per cent annual growth rate. Science and Public Policy, 1(6):122– 125, June 1974. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:AEF

[Ano75a] Anonymous. Adult education in the Federal Republic of Germany. Science and Public Policy, 2(4): 163–165, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:AC

[Ano75b] Anonymous. Agriculture in China. Science and Public Policy, 2(9):387–389, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:ATP

[Ano75c] Anonymous. Agriculture threatened by progress. Science and Public Policy, 2(9):392–394, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:AAS

[Ano75d] Anonymous. ASTEC: The Australian science Council. Science and Public Policy, 2(4):158– 160, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:BAP

[Ano75e] Anonymous. Backlash in São Paulo. Science and Public Policy, 2(3):131, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:BTL

[Ano75f] Anonymous. Better towns with less traffic. Science and Public Policy, 2(7):321–323, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:BNA

[Ano75g] Anonymous. Brazil’s nuclear agreement with West Germany. Science and Public Policy, 2(11): 502–503, November 1975. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:CUE

[Ano75h] Anonymous. Call for US energy policy. Science and Public Policy, 2(3):134–135, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:CWR

[Ano75i] Anonymous. China 1: Woman as a revolutionary force. Science and Public Policy, 2(6):252– 255, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:CWB

[Ano75j] Anonymous. China 2: Working as a barefoot doctor. Science and Public Policy, 2(6):256– 257, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:COS

[Ano75k] Anonymous. China 3: Observations on science. Science and Public Policy, 2(6):257–260, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:CRS

[Ano75l] Anonymous. Community research at the service of Europe. Science and Public Policy, 2(12):545–547, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:COR

[Ano75m] Anonymous. Curiosity oriented research. Science and Public Policy, 2(7):320, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:DDT

[Ano75n] Anonymous. Draft declaration of Tokyo. Science and Public Policy, 2(10):457, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:ECA

[Ano75o] Anonymous. Energy crisis in agriculture. Science and Public Policy, 2(7):300–303, July 1975. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:ESE

[Ano75p] Anonymous. Energy situation in the EEC. Science and Public Policy, 2(11):503–505, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:EMR

[Ano75q] Anonymous. Environment ministers’ resolutions. Science and Public Policy, 2(1):21–23, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:ECP

[Ano75r] Anonymous. Ethical code of Polish scientific workers. Science and Public Policy, 2(8):352– 353, August 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFa

[Ano75s] Anonymous. Facts in figures. Science and Public Policy, 2(1):42– 48, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFb

[Ano75t] Anonymous. Facts in figures.

Science and Public Policy, 2(6): 251, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFc

[Ano75u] Anonymous. Facts in figures.

Science and Public Policy, 2(6): 260, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFd

[Ano75v] Anonymous. Facts in figures. Science and Public Policy, 2(6):281– 287, June 1975. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFe

[Ano75w] Anonymous. Facts in figures. Science and Public Policy, 2(7):330– 336, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFf

[Ano75x] Anonymous. Facts in figures. Science and Public Policy, 2(8):365– 368, August 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFg

[Ano75y] Anonymous. Facts in figures. Science and Public Policy, 2(10):469– 472, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFh

[Ano75z] Anonymous. Facts in figures. Science and Public Policy, 2(11):514– 520, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FFi

[Ano75-27] Anonymous. Facts in figures. Science and Public Policy, 2(12):558– 566, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:FT

[Ano75-28] Anonymous. Fish for tomorrow. Science and Public Policy, 2(11):

509–510, November 1975. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:GPS

[Ano75-29] Anonymous. A general plan of the Swedish research councils. Science and Public Policy, 2(3):142– 144, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:GRO

[Ano75-30] Anonymous. Government role in offshore engineering. Science and Public Policy, 2(1):24, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:I

[Ano75-31] Anonymous. Index 1975. Science and Public Policy, 2(1):i– ??, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:MP

[Ano75-32] Anonymous. A matter of people.

Science and Public Policy, 2(10): 429–431, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:NRP

[Ano75-33] Anonymous. Neglected research priorities. Science and Public Policy, 2(7):323–325, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Na

[Ano75-34] Anonymous. News. Science and Public Policy, 2(1):??, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nb

[Ano75-35] Anonymous. News. Science and Public Policy, 2(2):50–55, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nc

[Ano75-36] Anonymous. News. Science and Public Policy, 2(3):99–109, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nd

[Ano75-37] Anonymous. News. Science and Public Policy, 2(4):146–151, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Ne

[Ano75-38] Anonymous. News. Science and Public Policy, 2(5):186–193, May 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nf

[Ano75-39] Anonymous. News. Science and Public Policy, 2(6):243–251, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Ng

[Ano75-40] Anonymous. News. Science and Public Policy, 2(7):290–295, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nh

[Ano75-41] Anonymous. News. Science and Public Policy, 2(8):338–345, August 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Ni

[Ano75-42] Anonymous. News. Science and Public Policy, 2(9):370–380, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nj

[Ano75-43] Anonymous. News. Science and Public Policy, 2(10):410–415, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nk

[Ano75-44] Anonymous. News. Science and Public Policy, 2(11):474–478, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Nl

[Ano75-45] Anonymous. News. Science and Public Policy, 2(12):522–529, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:NGM

[Ano75-46] Anonymous. Normal growth from a modified dynamic world model. Science and Public Policy, 2(3): 119–121, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:NFa

[Ano75-47] Anonymous. The nurturing forest. Science and Public Policy, 2 (9):395, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:OFG

[Ano75-48] Anonymous. Oilfield flared gas. Science and Public Policy, 2(2): 95–96, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:OPP

[Ano75-49] Anonymous. Organizational problems of the Polish Academy of Sciences. Science and Public Policy, 2(4):161–163, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:PPH

[Ano75-50] Anonymous. The place of public health in the economy of Africa. Science and Public Policy, 2(11): 493–497, November 1975. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:PUS

[Ano75-51] Anonymous. Planning for a uranium shortage. Science and Public Policy, 2(1):19–21, January 1975.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pa

[Ano75-52] Anonymous. Publications. Science and Public Policy, 2(1):35– 36, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pb

[Ano75-53] Anonymous. Publications. Science and Public Policy, 2(2):91– 94, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pc

[Ano75-54] Anonymous. Publications. Science and Public Policy, 2(3):139– 141, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pd

[Ano75-55] Anonymous. Publications. Science and Public Policy, 2(4):183– 184, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pe

[Ano75-56] Anonymous. Publications. Science and Public Policy, 2(6):274– 281, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pf

[Ano75-57] Anonymous. Publications. Science and Public Policy, 2(7):326– 329, July 1975. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pg

[Ano75-58] Anonymous. Publications. Science and Public Policy, 2(8):363– 364, August 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Ph

[Ano75-59] Anonymous. Publications. Science and Public Policy, 2(9):396– 402, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pi

[Ano75-60] Anonymous. Publications. Science and Public Policy, 2(10):461– 468, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pj

[Ano75-61] Anonymous. Publications. Science and Public Policy, 2(11):510– 513, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:Pk

[Ano75-62] Anonymous. Publications. Science and Public Policy, 2(12):554– 557, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:RDT

[Ano75-63] Anonymous. R and D trends in the OECD. Science and Public Policy, 2(12):538–542, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:RMP

[Ano75-64] Anonymous. Raw materials: a policy for the European Community. Science and Public Policy, 2(12):543–544, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:RNR

[Ano75-65] Anonymous. The role of the national research Council. Science and Public Policy, 2(9):390–391, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SIF

[Ano75-66] Anonymous. Saving India from floods and drought. Science and Public Policy, 2(4):176–177, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:STB

[Ano75-67] Anonymous. Science and techology in Brasil: a plan for national development. Science and Public Policy, 2(1):25, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SVP

[Ano75-68] Anonymous. Science in the VI-Ith plan. Science and Public Policy, 2(3):124–130, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SV

[Ano75-69] Anonymous. Science in Vojvodina. Science and Public Policy, 2(2):65–66, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SPD

[Ano75-70] Anonymous. Science policies in developing countries. Science and Public Policy, 2(2):56– 62, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SPN

[Ano75-71] Anonymous. Science policy in The Netherlands: a new approach: from our national correspondent. Science and Public Policy, 2(3): 110–118, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:STN

[Ano75-72] Anonymous. Science, technology and the new international economic order. Science and Public Policy, 2(2):66–74, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SRR

[Ano75-73] Anonymous. Scientific research in Romania. Science and Public Policy, 2(2):64–65, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SAG

[Ano75-74] Anonymous. Sickness absence in great Britain. Science and Public Policy, 2(7):310–311, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SSPa

[Ano75-75] Anonymous. Sociology of science priorities. Science and Public Policy, 2(2):62–63, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SAP

[Ano75-76] Anonymous. Soviet agriculture:

The past ten years. Science and Public Policy, 2(6):273, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SAa

[Ano75-77] Anonymous. SPF activities. Science and Public Policy, 2(6): 288, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SAb

[Ano75-78] Anonymous. SPF activities. Science and Public Policy, 2(11):508, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SAc

[Ano75-79] Anonymous. SPF activities. Science and Public Policy, 2(12):567– 568, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SSPb

[Ano75-80] Anonymous. STR and social progress. Science and Public Policy, 2(4):152–157, April 1975. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SRU

[Ano75-81] Anonymous. Student research in the USSR. Science and Public Policy, 2(3):133, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:SAS

[Ano75-82] Anonymous. Sweden: Allocations of science budget. Science and Public Policy, 2(3): 130, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:TTC

[Ano75-83] Anonymous. Technology transfer in the classics. Science and Public Policy, 2(2):63, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:TMM

[Ano75-84] Anonymous. Trends in medical manpower. Science and Public Policy, 2(6):261–264, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:UMR

[Ano75-85] Anonymous. UK Medical Research Council: Development of scientific policy. Science and Public Policy, 2(11):506, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:UIS

[Ano75-86] Anonymous. USA institute of the Soviet Academy of Sciences. Science and Public Policy, 2(3):132– 133, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:WB

[Ano75-87] Anonymous. War and bread.

Science and Public Policy, 2(6): 242, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:WYL

[Ano75-88] Anonymous. The way you lived and died. Science and Public Policy, 2(9):403–408, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:WDW

[Ano75-89] Anonymous. What do we gain from space flight? Science and Public Policy, 2(10):448–449, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:WWL

[Ano75-90] Anonymous. What will Lomé lead to? Science and Public Policy, 2(10):458–460, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:WR

[Ano75-91] Anonymous. Who is responsible?

Science and Public Policy, 2(12): 553–554, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1975:WF

[Ano75-92] Anonymous. World food. Science and Public Policy, 2(9):394, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:BFH

[Ano76a] Anonymous. Benjamin Franklin House. Science and Public Policy, 3(3):198–201, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:CD

[Ano76b] Anonymous. The Caribbean dilemma. Science and Public Policy, 3(2):152–159, April 1976. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:CVP

[Ano76c] Anonymous. Centralism vs pluralism in Australian science policy: The Report on Science to the Royal Commission on Australian Government Administration. Science and Public Policy, 3(2):148– 152, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:CEG

[Ano76d] Anonymous. Commonwealth experts group. Science and Public Policy, 3(4):339–341, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:CFT

[Ano76e] Anonymous. Commonwealth fund for technical cooperation. Science and Public Policy, 3(1):23– 24, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:CPB

[Ano76f] Anonymous. Control of pollution in Brasil. Science and Public Policy, 3(3):243–244, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:DSS

[Ano76g] Anonymous. Disciplines of social sciences. Science and Public Policy, 3(1):78–79, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:EPD

[Ano76h] Anonymous. Employee participation: a democratic necessity. Science and Public Policy, 3(3):247– 251, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:ERN

[Ano76i] Anonymous. Energy research in The Netherlands. Science and Public Policy, 3(1):57–59, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:E

[Ano76j] Anonymous. Europe +30. Science and Public Policy, 3(2):160– 162, April 1976. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:FFa

[Ano76k] Anonymous. Facts in figures. Science and Public Policy, 3(1): 11, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:FFb

[Ano76l] Anonymous. Facts in figures. Science and Public Policy, 3(2):173– 179, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:FFc

[Ano76m] Anonymous. Facts in figures. Science and Public Policy, 3(3):285– 288, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:FFd

[Ano76n] Anonymous. Facts in figures. Science and Public Policy, 3(4):380– 384, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:IRD

[Ano76o] Anonymous. Increase in R&D funding forecast. Science and Public Policy, 3(2):113, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:I

[Ano76p] Anonymous. Index. Science and Public Policy, 3(2):180, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:TEU

[Ano76q] Anonymous. Is there enough uranium. Science and Public Policy, 3(4):350–355, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:TMP

[Ano76r] Anonymous. It’s time to make paper in the tropics. Science and Public Policy, 3(2):108–112, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:LTG

[Ano76s] Anonymous. Long-term global costs of nuclear weapons. Science and Public Policy, 3(1):46– 57, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:NMP

[Ano76t] Anonymous. Netherland minister presents new-style science budget. Science and Public Policy, 3 (1):21–22, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Na

[Ano76u] Anonymous. News. Science and Public Policy, 3(1):2–11, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Nb

[Ano76v] Anonymous. News. Science and Public Policy, 3(2):90–102, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Nc

[Ano76w] Anonymous. News. Science and Public Policy, 3(3):182–197, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Nd

[Ano76x] Anonymous. News. Science and Public Policy, 3(4):290–295, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:ODA

[Ano76y] Anonymous. Obstacles to the development of arid and semi-arid zones. Science and Public Policy, 3(3):224–237, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:PUG

[Ano76z] Anonymous. Pregnancy and the unmarried girl. Science and Public Policy, 3(4):347–350, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Pa

[Ano76-27] Anonymous. Publications. Science and Public Policy, 3(1):80– 88, January 1976. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Pb

[Ano76-28] Anonymous. Publications. Science and Public Policy, 3(2):165– 172, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Pc

[Ano76-29] Anonymous. Publications. Science and Public Policy, 3(3):269– 281, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:Pd

[Ano76-30] Anonymous. Publications. Science and Public Policy, 3(4):373– 379, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:QWL

[Ano76-31] Anonymous. The quality of working life. Science and Public Policy, 3(1):28–30, January 1976. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:RBC

[Ano76-32] Anonymous. Raising biological ceilings. Science and Public Policy, 3(1):79–80, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:RUE

[Ano76-33] Anonymous. Rational use of energy. Science and Public Policy, 3 (3):246–247, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:STF

[Ano76-34] Anonymous. Science task force to assist Royal Commission. Science and Public Policy, 3(1):31, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:SAT

[Ano76-35] Anonymous. Seminar on alternative technology. Science and Public Policy, 3(1):19–20, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:SES

[Ano76-36] Anonymous. Significance of European security act. Science and Public Policy, 3(4):342–347, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:SAa

[Ano76-37] Anonymous. SPF activities. Science and Public Policy, 3(1):77, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:SAb

[Ano76-38] Anonymous. SPF activities. Science and Public Policy, 3(2): 162, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:SAc

[Ano76-39] Anonymous. SPF activities. Science and Public Policy, 3(4):379, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:TP

[Ano76-40] Anonymous. Teachers and politics. Science and Public Policy, 3 (1):27–28, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:UIO

[Ano76-41] Anonymous. USSR industrial output. Science and Public Policy, 3(4):368–369, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:WRN

[Ano76-42] Anonymous. Where research is needed. Science and Public Policy, 3(2):163–164, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1976:WSO

[Ano76-43] Anonymous. WMO statement on ozone. Science and Public Policy, 3(3):237–243, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:AST

[Ano77a] Anonymous. Application of science and technology in the Arab world. Science and Public Policy, 4(4):355–358, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:BCA

[Ano77b] Anonymous. Benefit-cost analysis. Science and Public Policy, 4 (2):184–187, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:BNP

[Ano77c] Anonymous. Brazil’s nuclear power programme. Science and Public Policy, 4(1):85–87, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:CNC

[Ano77d] Anonymous. Capital needs: can we afford the future? Science and Public Policy, 4(1):53– 54, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:CIS

[Ano77e] Anonymous. Crisis in Indian science. Science and Public Policy, 4(6):583–584, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:EIS

[Ano77f] Anonymous. Educational institutions and scientific disciplines. Science and Public Policy, 4(1): 64–68, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:EAD

[Ano77g] Anonymous. Energy for agriculture in the developing countries. Science and Public Policy, 4(2):

168–179, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:FFa

[Ano77h] Anonymous. Facts in figures. Science and Public Policy, 4(1):100– 104, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:FFb

[Ano77i] Anonymous. Facts in figures. Science and Public Policy, 4(6):578– 581, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:FTN

[Ano77j] Anonymous. Food taboos and nutrition. Science and Public Policy, 4(1):94–96, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:HHR

[Ano77k] Anonymous. Health as a human right. Science and Public Policy, 4(2):188–191, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:HHG

[Ano77l] Anonymous. The health hazards of not going nuclear. Science and Public Policy, 4(3):291– 296, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:HRN

[Ano77m] Anonymous. Humanities research in Netherlands science budget.

Science and Public Policy, 4(3): 219–224, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:IA

[Ano77n] Anonymous. Impressive averages. Science and Public Policy, 4(1): 92–93, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:I

[Ano77o] Anonymous. Index 1977. Science and Public Policy, 4(6):??, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:WMI

[Ano77p] Anonymous. ‘… What’s more, it burns money and is guaranteed to lay a mile of unemployed an hour’. Science and Public Policy, 4(6):538–539, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:LLR

[Ano77q] Anonymous. Low-level radioactive waste. Science and Public Policy, 4(3):239–242, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:MRC

[Ano77r] Anonymous. The medical research Council and inflation. Science and Public Policy, 4(1):55– 59, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:MMM

[Ano77s] Anonymous. More meat, more waste. Science and Public Policy, 4(1):93–94, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Na

[Ano77t] Anonymous. News. Science and Public Policy, 4(2):107–118, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Nb

[Ano77u] Anonymous. News. Science and Public Policy, 4(3):210–218, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Nc

[Ano77v] Anonymous. News. Science and Public Policy, 4(4):341–345, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Nd

[Ano77w] Anonymous. News. Science and Public Policy, 4(5):394–396, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Ne

[Ano77x] Anonymous. News. Science and Public Policy, 4(6):490–504, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:NCT

[Ano77y] Anonymous. Nothing can be taken as given. Science and Public Policy, 4(2):200–201, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:PMA

[Ano77z] Anonymous. Policy making in Arab countries. Science and Public Policy, 4(1):21–34, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:PDa

[Ano77-27] Anonymous. Press digest. Science and Public Policy, 4(4):345– 350, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:PDb

[Ano77-28] Anonymous. Press digest. Science and Public Policy, 4(5):454– 458, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:PAS

[Ano77-29] Anonymous. Public attitudes to science. Science and Public Policy, 4(1):38–49, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:PIS

[Ano77-30] Anonymous. Public involvement in science and technology. Science and Public Policy, 4(1):49– 51, February 1977. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Pa

[Ano77-31] Anonymous. Publications. Science and Public Policy, 4(1):96– 100, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Pb

[Ano77-32] Anonymous. Publications. Science and Public Policy, 4(3):297– 303, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Pc

[Ano77-33] Anonymous. Publications. Science and Public Policy, 4(5):488, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:Pd

[Ano77-34] Anonymous. Publications. Science and Public Policy, 4(6):573– 578, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:SRD

[Ano77-35] Anonymous. Science and regulatory decision: chemicals in the environment. Science and Public Policy, 4(2):163–166, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:SAa

[Ano77-36] Anonymous. SPF activities. Science and Public Policy, 4(3):

304, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:SAb

[Ano77-37] Anonymous. SPF activities. Science and Public Policy, 4(6):582, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:SE

[Ano77-38] Anonymous. Strategies for Europe. Science and Public Policy, 4 (2):106, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:TAB

[Ano77-39] Anonymous. Technology assessment in business and government. Science and Public Policy, 4(3): 229–231, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:UP

[Ano77-40] Anonymous. Uranium prospect. Science and Public Policy, 4(2): 167–168, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1977:WBA

[Ano77-41] Anonymous. The world bank and aid for health projects. Science and Public Policy, 4(3):244– 249, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:AH

[Ano78a] Anonymous. Aid for health. Science and Public Policy, 5(2):147– 153, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:CCS

[Ano78b] Anonymous. Canada as a conserver society. Science and Public Policy, 5(1):66–68, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:EE

[Ano78c] Anonymous. East Europe. Science and Public Policy, 5(2): 160, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:EIU

[Ano78d] Anonymous. Energy investment in the UK. Science and Public Policy, 5(3):208–217, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:ETY

[Ano78e] Anonymous. Europe tackles youth unemployment. Science and Public Policy, 5(2):135–138, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:FFa

[Ano78f] Anonymous. Facts in figures.

Science and Public Policy, 5(4): 320, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:FFb

[Ano78g] Anonymous. Facts in figures. Science and Public Policy, 5(6):479– 480, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:GMO

[Ano78h] Anonymous. Generating methane from organic wastes. Science and Public Policy, 5(6):412–416, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:I

[Ano78i] Anonymous. Index 1978. Science and Public Policy, 5(6):481–483, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:IIS

[Ano78j] Anonymous. International indicators of science and technology. Science and Public Policy, 5(3): 195–205, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:IT

[Ano78k] Anonymous. International tourism. Science and Public Policy, 5(2): 154–158, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:NPM

[Ano78l] Anonymous. A national policymaking body for science and technology in Jordan. Science and Public Policy, 5(3):218–231, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:N

[Ano78m] Anonymous. News. Science and Public Policy, 5(1):12–26, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:NRP

[Ano78n] Anonymous. Norwegian R&D policy. Science and Public Policy, 5 (4):247–250, August 1978. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:NEU

[Ano78o] Anonymous. Nuclear energy in the USSR. Science and Public Policy, 5(2):122–123, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:NSQ

[Ano78p] Anonymous. Nuclear science and questions shaping its future. Science and Public Policy, 5(1):63– 64, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:OCP

[Ano78q] Anonymous. Occupational cancer: prevention through public information. Science and Public Policy, 5(1):43–44, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:POH

[Ano78r] Anonymous. Policy for occupational health. Science and Public Policy, 5(1):??, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:PPI

[Ano78s] Anonymous. Poor performance by Irish services sector. Science and Public Policy, 5(4):251– 253, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:Pa

[Ano78t] Anonymous. Publications. Science and Public Policy, 5(3):239– 240, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:Pb

[Ano78u] Anonymous. Publications. Science and Public Policy, 5(4):318– 320, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:RCA

[Ano78v] Anonymous. Real cost of the arms race. Science and Public Policy, 5(1):52, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:RFT

[Ano78w] Anonymous. Research on the future and technology assessment. Science and Public Policy, 5(6): 471–474, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:SAa

[Ano78x] Anonymous. SPF activities. Science and Public Policy, 5(1):79– 80, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:SAb

[Ano78y] Anonymous. SPF activities. Science and Public Policy, 5(5):400, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:SDM

[Ano78z] Anonymous. Strategy for the development and management of deserts. Science and Public Policy, 5(1):56–62, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:NSP

[Ano78-27] Anonymous. The Netherlands: 1.

science policy. Science and Public Policy, 5(4):242–246, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:NPP

[Ano78-28] Anonymous. The Netherlands: 2.

planning policy. Science and Public Policy, 5(4):246–247, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:TUS

[Ano78-29] Anonymous. Trade unions in Soviet society. Science and Public Policy, 5(6):475–478, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:UFR

[Ano78-30] Anonymous. US Federal R&D growth. Science and Public Policy, 5(2):158–159, April 1978. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:UEE

[Ano78-31] Anonymous. Use of energy in European agriculture. Science and Public Policy, 5(2):110–122, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:UE

[Ano78-32] Anonymous. USSR economy in 1977. Science and Public Policy, 5 (3):236, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:VNF

[Ano78-33] Anonymous. Vision of the near future unless. … Science and Public Policy, 5(1):47, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1978:YUO

[Ano78-34] Anonymous. Youth unemployment in the OEeD. Science and Public Policy, 5(3):237–239, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:AH

[Ano79a] Anonymous. Ad hoc. Science and Public Policy, 6(3):200– 202, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:BDG

[Ano79b] Anonymous. Benefits and dangers of genetic engineering. Science and Public Policy, 6(3):150– 154, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:CPU

[Ano79c] Anonymous. Comparative performance of US technology. Science and Public Policy, 6(3):154– 162, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:DSP

[Ano79d] Anonymous. Dutch science policy:

Highlights from RAWB’s annual report. Science and Public Policy, 6(6):447–451, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:EPP

[Ano79e] Anonymous. Environmental policies and prospects for the 1980s. Science and Public Policy, 6(5): 345–359, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:ECP

[Ano79f] Anonymous. European competition policy. Science and Public Policy, 6(6):451–455, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:FFa

[Ano79g] Anonymous. Facts in figures. Science and Public Policy, 6(3):196– 199, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:FFb

[Ano79h] Anonymous. facts in figures. Science and Public Policy, 6(6):455, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:I

[Ano79i] Anonymous. Index 1979. Science and Public Policy, 6(6):462–464, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:NBS

[Ano79j] Anonymous. National budgeting of science and technology activities. Science and Public Policy, 6 (2):120–133, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:PAa

[Ano79k] Anonymous. The prospects for Africa. Science and Public Policy, 6(6):420–423, December 1979.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:Pa

[Ano79l] Anonymous. Publications. Science and Public Policy, 6(1):70– 72, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:Pb

[Ano79m] Anonymous. Publications. Science and Public Policy, 6(3):203– 207, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:Pc

[Ano79n] Anonymous. Publications. Science and Public Policy, 6(5):382– 384, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:Pd

[Ano79o] Anonymous. Publications. Science and Public Policy, 6(6):461, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:RRS

[Ano79p] Anonymous. Research related to social planning in Norway. Science and Public Policy, 6(2): 133, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:SCE

[Ano79q] Anonymous. Sociocultural consequences of ecosystem transformation. Science and Public Policy, 6 (3):166–170, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:SPG

[Ano79r] Anonymous. Survey of S and T policies of the German democratic Republic. Science and Public Policy, 6(2):82–85, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:TYR

[Ano79s] Anonymous. Ten years of RAWB: From our special correspondent — The Hague. Science and Public Policy, 6(2):112–120, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Anonymous:1979:TFE

[Ano79t] Anonymous. Tropical forest ecosystems. Science and Public Policy, 6(3):164–166, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Archer:1978:EPA

[Arc78]  Bruce Archer.  Epistemology

and the practical arts. Science and Public Policy, 5(4):304– 311, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Armitage:1975:PCC

[Arm75] E. Armitage. Patents conference — the current state of play. Science and Public Policy, 2(1):13– 18, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Al-Saadi:1978:PO

[AS78] Ahmed Al-Saadi. Prospects for oil. Science and Public Policy, 5 (3):205–208, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Akerele:1977:NRM

[ATM77] O. Akerele, I. Tabibzadeh, and J. McGilvray. New role for medical missionaries. Science and Public Policy, 4(3):267–275, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aurum:1974:AHa

[Aur74a] Aurum. ad hoc. Science and Public Policy, 1(10):309–310, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aurum:1974:AHb

[Aur74b] Aurum.  Ad hoc.  Science

and Public Policy, 1(11):342–343, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aurum:1974:AHc

[Aur74c] Aurum. Ad hoc. Science and Public Policy, 1(12):388–389, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aurum:1974:A

[Aur74d] Aurum.  Adhoc.  Science

and Public Policy, 1(9):256–257, September 1974. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Avebury:1975:SLG

[Ave75] Avebury. Social limits to growth. Science and Public Policy, 2(3): 98, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Aziz:1976:RDC

[Azi76] Sartaj Aziz. Rural development: The Chinese approach. Science and Public Policy, 3(1):12– 18, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bard:1975:PLR

[Bar75a] Dr Basil Bard. Patent law reform. Science and Public Policy, 2(7): 317–319, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Barzalatto:1975:SRS

[Bar75b] J. Barzalatto. Self- reliance 2: Self- reliance and international collaboration in science and technology. Science and Public Policy, 2(10):424–429, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bard:1977:DL

[Bar77] Dr B. Bard. The disjunctive leap. Science and Public Policy, 4(4): 363–364, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Burton-Bradley:1977:PIP

[BB77]  B. G. Burton-Bradley.  Psy-

chosocial implications of Papua New Guinean independence. Science and Public Policy, 4(2):195– 200, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bakker:1977:POS

[BC77] C. J. Bakker and K. H. Chang.

Publication output in the sciences: a 20-country survey. Science and Public Policy, 4(6):563–565, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1979:SEA

[BC79a] M. R. Brett-Crowther. Strategies for Europe and America. Science and Public Policy, 6(6):456–461, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brett-Crowther:1979:FFA

[BC79b] Michael Brett-Crowther. The force field around science. Science and Public Policy, 6(5):379– 382, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Beaton:1976:DFN

[Bea76] G. H. Beaton. Defining a food and nutrition policy. Science and Public Policy, 3(4):308–321, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bekoe:1979:SSS

[Bek79] D. A. Bekoe. Statement of the Singapore symposium. Science and Public Policy, 6(4):213– 216, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bentley:1974:RA

[Ben74] Barbara Bentley. Rights of Aborigines. Science and Public Policy, 1(10):287–290, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Berrie:1978:TTN

[Ber78a] T. W. Berrie. Technology transfer: a new approach to training. Science and Public Policy, 5(5):345– 348, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Berrie:1978:TPE

[Ber78b] T. W. Berrie. Transfer of project engineering and design skills to developing countries. Science and Public Policy, 5(4):284–290, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Berrie:1978:TTA

[Ber78c] Tw Berrie. Technological transfer: 1. appropriateness for developing countries. Science and Public Policy, 5(6):444–448, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Backstrand:1975:SCQ

[BI75] Göran Bäckstrand and Lars Ingelstam. Sweden as a case in the quest for appropriate development. Science and Public Policy, 2(10):431–441, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Birman:1976:FS

[Bir76] Alexander Birman. Financing science. Science and Public Policy, 3(4):367–368, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Berrie:1978:TTT

[BL78] T. W. Berrie and D. Leslie. Technological transfer: 2. through developing small businesses. Science and Public Policy, 5(6):448–451, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Black:1979:BBR

[Bla79] Guy Black. Big is better read. Science and Public Policy, 6(4):294– 296, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boeker:1977:PIS

[Boe77] Egbert Boeker. Public information on science and technology: the Dutch case. Science and Public Policy, 4(6):558–562, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Boffey:1976:RIR

[Bof76] Philip M. Boffey. Rasmussen issues revised odds on a nuclear catastrophe. Science and Public Policy, 3(2):146–148, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bognar:1977:ONS

[Bog77] József Bognar. Outlines of a new system of international economic relations. Science and Public Policy, 4(6):523–533, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bojko:1979:NIE

[Boj79] Béla Bojko. New international economic order and flow of science and technology. Science and Public Policy, 6(6):424–432, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borgstrom:1974:FPD

[Bor74] Georg Borgstrom. The food– population dilemma. Science and Public Policy, 1(12):406–412, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Borgstrom:1975:PT

[Bor75] Georg Borgstrom. The price of a tractor. Science and Public Policy, 2(8):345–351, August 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Bowonder:1979:TTS

[Bow79] B. Bowonder. Technology transfer: a strategy for the 1980s. Science and Public Policy, 6(4):266– 275, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braun:1977:CTI

[Bra77a] Ernest Braun. Control of technology in interest of health. Science and Public Policy, 4(6):540–544, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Braun:1977:TAR

[Bra77b] Ernest Braun. Technology assessment and the role of the universities. Science and Public Policy, 4 (3):224–229, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brezhnev:1976:SS

[Bre76] Leonid Brezhnev. Socialism and science. Science and Public Policy, 3(1):24–26, January 1976. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brookes:1976:EAN

[Bro76] L. G. Brookes. Energy accounting and nuclear power. Science and Public Policy, 3(1):60– 64, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Brown:1978:LSE

[Bro78] Lord (Wilfred) Brown. Learning from the Swedish experience. Science and Public Policy, 5(3):233– 235, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Budowski:1975:SI

[Bud75] Gerardo Budowski. Scientific imperialism. Science and Public Policy, 2(8):354–360, August 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Colombo:1974:RDP

[CA74] Umberto Colombo and Gabriel Arilhon. R and D policy for materials resources. Science and Public Policy, 1(11):330–334, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Carter:1974:FD

[Car74] Charles F. Carter. Future darkness 2. Science and Public Policy, 1(6):99–100, June 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Carter:1976:WFC

[Car76] James P. Carter. World food crisis. Science and Public Policy, 3(4):296–308, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Carter:1979:SE

[Car79] Sir Charles Carter. Strategies for Europe. Science and Public Policy, 6(2):142–144, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cassirer:1978:CT

[Cas78] Henry R. Cassirer. Challenge of the telephone. Science and Public Policy, 5(4):311–315, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cepede:1976:SN

[Cép76] Michel Cépède. Sociology and nutrition. Science and Public Policy, 3(2):114–119, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cherns:1977:SDA

[Che77] Albert Cherns. Social development and the assessment of technology. Science and Public Policy, 4(4):350–355, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cherns:1978:SST

[Che78] A. B. Cherns. Sociology of scientific and technological policy. Science and Public Policy, 5(1):68– 79, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clarke:1978:GE

[Cla78] Arthur C. Clarke. God and Einstein. Science and Public Policy, 5(5):399–400, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clemison:1974:SSS

[Cle74] John Clemison. Social stratification in science. Science and Public Policy, 1(9):258–259, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clemison:1975:CS

[Cle75] John Clemison. Communication in science. Science and Public Policy, 2(1):34–35, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Clerk:1979:NEP

[Cle79] Norman Clerk. Nigerian experience with policy research. Science and Public Policy, 6(4):229– 237, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Coles:1975:RFP

[Col75] James Stacy Coles. Report of foundation President. Science and Public Policy, 2(5):234–237, May 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Colombo:1977:IIE

[Col77] Umberto Colombo. Industrial innovation in Europe. Science and Public Policy, 4(4):326–339, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Collingridge:1979:ETC

[Col79] D. G. Collingridge. The entrenchment of technology: the case of lead petrol additives. Science and Public Policy, 6(5):332– 338, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cooper:1976:HRP

[Coo76] Michael H. Cooper. Health resources: planning and priorities. Science and Public Policy, 3(1): 36–40, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Cooper:1977:CCM

[Coo77] Paul Cooper. The case for commercial money. Science and Public Policy, 4(1):51–53, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Correspondent:1975:SPN

[Cor75] National Correspondent. Science policy in The Netherlands: a new approach. Science and Public Policy, 2(4):166–176, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Councils:1977:CSU

[Cou77] Research Councils. Civil science (United Kingdom). Science and Public Policy, 4(3): 283, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Craig:1976:ECP

[Cra76] Dennis Craig. Erosion control provides fresh fish. Science and Public Policy, 3(4):321–322, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Crozier:1978:SEI

[Cro78] Michel Crozier. Structural evolution in industrialized societies.

Science and Public Policy, 5(5): 385–393, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Darling:1975:FEM

[Dar75] Frank Fraser Darling. Forestry, the environment and man’s needs. Science and Public Policy, 2(7): 304–309, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

daCosta:1978:IRS

[dC78a] Joao Frank da Costa. II. reflections by the secretary-general. Science and Public Policy, 5(1):5– 11, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

diCastri:1978:PII

[dC78b] Francesco di Castri. Planning international interdisciplinary research. Science and Public Policy, 5(4):254–266, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Darin-Drabkin:1974:DCU

[DD74] Haim Darin-Drabkin. Development and change in the USSR. Science and Public Policy, 1(10): 302–308, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Darin-Drabkin:1975:PSV

[DD75] Dr Haim Darin-Drabkin. Is a Palestinian state viable? Science and Public Policy, 2(10):450– 456, October 1975. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Dedijer:1975:TSI

[Ded75] Stevan Dedijer. Technology, science and industry in developing countries focus: India and China. Science and Public Policy, 2(5): 238–240, May 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dedijer:1978:IDL

[Ded78] S. Dedijer. An intelligence doctrine for the LDCs. Science and Public Policy, 5(5):333–345, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

deHemptinne:1975:SPN

[dH75] Y. de Hemptinne. Science Policy in the Nordic countries. Science and Public Policy, 2(7):296– 300, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dick:1979:QAB

[Dic79] K. J. Dick. Qu’Appelle Basin Study: public involvement program. Science and Public Policy, 6(3):178–184, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dobrov:1974:SPS

[Dob74] G. M. Dobrov. Science policy in the Soviet union. Science and Public Policy, 1(12):402–405, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dobrov:1978:INI

[DRR78] G. M. Dobrov, R. H. Randolph, and W. D. Rauch. Informational networks for international team research. Science and Public Policy, 5(5):322–333, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Drury:1974:SSF

[Dru74] C. M. Drury. Social science funding in Canada. Science and Public Policy, 1(12):??, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Price:1978:RNH

[dSP78] Derek de Solla Price. Revolution needs heroism. Science and Public Policy, 5(3):190–195, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dubos:1976:PD

[Dub76] Rene Dubos. Progressive degradation. Science and Public Policy, 3 (1):71–77, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dunning:1974:FME

[Dun74] John H. Dunning. The future of the multinational enterprise. Science and Public Policy, 1(9):233– 238, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Dyomin:1977:IB

[Dyo77] Valeri Dyomin. Institute of brains. Science and Public Policy, 4(5): 486–487, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Eisner:1979:HAE

[Eis79] H. S. Eisner. Hazard assessment:

experience and expectation. Science and Public Policy, 6(3):146– 150, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Eldin:1979:TE

[Eld79] Gérard Eldin. Tourism and the environment. Science and Public Policy, 6(1):60–61, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

El-Saadi:1976:AEF

[ES76] Ahmed El-Saadi. Arab–European fields of cooperation. Science and Public Policy, 3(2):119–121, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Evans:1976:NPA

[Eva76] Karl R. Evans. Nuclear power accident evaluation. Science and Public Policy, 3(1):69–70, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fairclough:1977:JEP

[Fai77] A. J. Fairclough. Japan environment policies. Science and Public Policy, 4(6):545–547, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fitzhugh:1979:SAA

[FB79] H. A. Fitzhugh and E. K. Byington. Systems approach to animal agriculture. Science and Public Policy, 6(1):43–49, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Feld:1975:NEF

[Fel75] B. T. Feld. Nuclear energy — fact vs myth. Science and Public Policy, 2(6):267–272, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fenn:1976:MSD

[Fen76]  M. G. Fenn.  Milk schemes

in developing countries. Science and Public Policy, 3(3):201– 207, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fishlock:1977:OSC

[Fis77] David Fishlock. ‘from our science correspondent’. Science and Public Policy, 4(3):275–280, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Flowers:1977:AES

[Flo77] Sir Brian Flowers. Approach of the European science foundation. Science and Public Policy, 4(4): 319–326, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Forrester:1976:LGR

[For76] Jay W. Forrester. Limits to growth revisited. Science and Public Policy, 3(2):143–146, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Freedman:1978:SFC

[Fre78] Ron Freedman. Science fragments culture. Science and Public Policy, 5(6):469–471, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1975:BS

[Fry75] Dr Gordon Fryers. The business of science. Science and Public Policy, 2(11):507–508, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1977:IVR

[Fry77a] Gordon Fryers. Industry views of the role of the Federal Government in industrial innovation. Science and Public Policy, 4(6): 572–573, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1977:FIR

[Fry77b] Gordon R. Fryers. Financing innovative risk: 1. Science and Public Policy, 4(2):126–132, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Fryers:1979:MP

[Fry79] Gordon R. Fryers. Methodology for planners. Science and Public Policy, 6(3):196, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gabor:1974:AQL

[Gab74] D. Gabor. Aeronautics and the quality of life. Science and Public Policy, 1(8):183–186, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Garfield:1975:TFS

[Gar75] Dr Eugene Garfield. Is there a future for the scientific journal? Science and Public Policy, 2(11):498– 501, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gaudry:1975:RA

[Gau75] Roger Gaudry. The road ahead. Science and Public Policy, 2(12): 530–538, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1977:PS

[GD77] Maurice Goldsmith and Daniel Danin. ‘popularization’ of science. Science and Public Policy, 4(5): 458–464, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Gellhorn:1977:WME

[Gel77] Alfred Gellhorn. WHO and medical ethics. Science and Public Policy, 4(6):550–552, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Giarini:1979:EVD

[Gia79] Orio Giarini. Economics, vulnerability and diminishing returns of technology. Science and Public Policy, 6(5):318–331, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1979:ISS

[GK79] Maurice Goldsmith and Alexander King. International Symposium on Science and Technology for Development (ISSTA): Singapore, 22–27 January 1979: Introduction. Science and Public Policy, 6(4):211–213, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Godber:1975:NPH

[God75] G. E. Godber. New perspective on health of Canadians. Science and Public Policy, 2(4):182– 183, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Godfrey:1978:MPC

[God78] Martin Godfrey. Migration of professionals from Commonwealth developing countries. Science and Public Policy, 5(5):367–375, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1974:PS

[Gol74] Maurice Goldsmith. ‘popularization’ of science. Science and Public Policy, 1(12):392–396, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1977:IIS

[Gol77a] Maurice Goldsmith. Industrial innovation: strategies and risks. Science and Public Policy, 4(2): 119, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1977:SPE

[Gol77b] Maurice Goldsmith. Science policy for Europe. Science and Public Policy, 4(4):306, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1978:EB

[Gol78a] Maurice Goldsmith. The evolution of a biologist. Science and Public Policy, 5(3):232–233, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1978:RRG

[Gol78b] Maurice Goldsmith. Reflections on the role of government in the dissemination of knowledge. Science and Public Policy, 5(1):46– 52, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1979:OWI

[Gol79a] Maurice Goldsmith. One world intellectual community. Science and Public Policy, 6(6):443–446, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goldsmith:1979:SA

[Gol79b] Maurice Goldsmith. Search for answers. Science and Public Policy, 6(4):210–211, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Graham:1978:SRL

[Gra78] Roger Frederic Graham. Selfrefinement: the last resource. Science and Public Policy, 5(4):316– 317, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Goeller:1975:ASW

[GW75] H. E. Goeller and Alvin M. Weinberg. The age of substitutability or What do we do when the mercury runs out? Science and Public Policy, 2(11):479–492, November 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Haefele:1978:ED

[Hae78] W. Haefele. On energy demand.

Science and Public Policy, 5(2): 93–105, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hague:1979:GRD

[Hag79a] The Hague. Government and R&D in industry. Science and Public Policy, 6(1):49–53, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hague:1979:STI

[Hag79b] The Hague. Scientific and technological information. Science and Public Policy, 6(3):163–164, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hamburger:1975:PBM

[Ham75] Jean Hamburger. Physicians, biologists and man’s future. Science and Public Policy, 2(1):29– 33, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hawthorne:1977:FI

[Haw77] Edward Hawthorne. Framework for innovation. Science and Public Policy, 4(2):120–125, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hayes:1977:NT

[Hay77] Denis Hayes. Nuclear terrorism. Science and Public Policy, 4(1): 82–85, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hughes-Evans:1979:TCD

[HE79] David Hughes-Evans. The technician’s contribution to development. Science and Public Policy, 6(4):237–241, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hill:1976:NPA

[Hil76] Sir John Hill. Nuclear power: achievements, problems and myths. Science and Public Policy, 3(4):355–366, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hammond:1976:MRF

[HLRS76] S. B. Hammond, J. A. Lane, A. Rogov, and R. Skjoeldebrand. Manpower requirements for future nuclear power. Science and Public Policy, 3(1):64–69, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hodgkins:1977:UEG

[Hod77] Paul Hodgkins. Unexpected effects of government intervention. Science and Public Policy, 4(2): 142–148, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hoelscher:1979:STD

[Hoe79a] H. E. Hoelscher. Science, technology and development policy: an analysis of constraints. Science and Public Policy, 6(3):170– 177, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hoelscher:1979:TST

[Hoe79b] H. E. Hoelscher. Transfer of science and technology to Lebanon. Science and Public Policy, 6(2): 91–106, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Hornig:1979:HST

[Hor79] Donald Hornig. Health of the scientific and ‘technical’ enterprise. Science and Public Policy, 6(2): 107–111, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Icerman:1977:PWD

[Ice77] L. Icerman. Perspectives on the worldwide debate and public opinion on nuclear power. Science and Public Policy, 4(5):465–482, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Izmerov:1974:PUE

[Izm74] N. F. Izmerov. Principles underlying the establishment of air quality standards in USSR. Science and Public Policy, 1(9):230–233, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jansky:1978:ISS

[Jan78] Donald M. Jansky. Information systems for a sustainable society. Science and Public Policy, 5(6): 454–456, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jeffreys:1976:MQG

[Jef76] Margot Jeffreys. Measuring the quality of G.P. care. Science and Public Policy, 3(1):40–46, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Johnston:1977:ESP

[Joh77] Ron Johnston. European social policy today and tomorrow. Science and Public Policy, 4(6):569– 570, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Jones:1977:FIR

[Jon77] Sydney Jones. Financing innovative risk: 2. Science and Public Policy, 4(2):132–134, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Khamis:1976:WCA

[KA76] S. H. Khamis and D. C. Alonzo.

1980 world census of agriculture. Science and Public Policy, 3(3): 214–224, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kanga:1977:ICM

[Kan77] N’Guessan Kanga. The Ivory Coast Ministry of Scientific Research. Science and Public Policy, 4(1):34–38, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kates:1976:IRP

[Kat76] Joseph Kates. Industrial R&D policies. Science and Public Policy, 3(3):244–246, June 1976. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Katz:1978:QQI

[Kat78] Shaul Katz. Quantity and quality in Israeli science. Science and Public Policy, 5(5):348–358, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kapferer:1978:FAP

[KB78] Jean-Noël Kapferer and Jean-Francois Boss. French attitudes to the popularization of science.

Science and Public Policy, 5(3): 162–180, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krummel:1977:RCA

[KD77] J. Krummel and W. Dritschilo. Resource costs of animal protein production. Science and Public Policy, 4(3):249–254, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kearton:1977:IBD

[Kea77] Lord Kearton. Increasing Britain’s development effectiveness. Science and Public Policy, 4(3):284–288, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kendrew:1974:SCT

[Ken74] Sir John Kendrew. Science and the change of time. Science and Public Policy, 1(10):278–285, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

King:1974:EO

[Kin74] Alexander King. The elite and the obsolete. Science and Public Policy, 1(12):367–375, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

King:1975:PDI

[Kin75] Alexander King. Pavane for a defunct institution. Science and Public Policy, 2(2):75–78, February 1975. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

King:1978:PEP

[Kin78] K. F. S. King. The political economy of pulp and paper. Science and Public Policy, 5(5):376– 385, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kotter:1974:DDP

[Kot74] Herbert R. Kotter. Deficiencies in development policy. Science and Public Policy, 1(11):318, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kroller:1978:DPD

[Krö78] Edgar Kröller. The debt problem of developing countries. Science and Public Policy, 5(4):295– 303, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Krebsbach:1978:ARD

[KS78] Camilla Krebsbach and Joachim Scharioth. Attitudes regarding different forms of energy production. Science and Public Policy, 5(6):407–412, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Kubin:1978:VSP

[Kub78] Jerzy Kubin. Value of science in Poland. Science and Public Policy, 5(2):124–135, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lal:1974:SCS

[Lal74] Dr Shivaji Lal. Social control and social responsibility. Science and Public Policy, 1(10):297– 302, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lall:1975:MDN

[Lal75] Sanjaya Lall. Multinationals and development: a new look. Science and Public Policy, 2(7):312– 316, July 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lartius:1977:RFD

[Lar77] S. Lartius. A room full of dukes. Science and Public Policy, 4(4): 364–365, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Laszlo:1975:RSG

[Las75] Ervin Laszlo. Responsibility of science in a global society. Science and Public Policy, 2(2):85– 90, February 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lavie:1978:CII

[Lav78] A. Lavie. Constraints on Israeli industrial innovation. Science and Public Policy, 5(4):280– 284, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lawrence:1979:CSE

[Law79] Peter Lawrence. Conception and standing of engineering work in Britain and Germany. Science and Public Policy, 6(1):53–59, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leger:1977:CAU

[Leg77] J. M. Leger. Challenge to African universities. Science and Public Policy, 4(1):80–82, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Leontief:1978:EPA

[Leo78] Wassily Leontief. Employment policies in the age of automation. Science and Public Policy, 5(6): 451–453, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Linder:1976:RTH

[Lin76] Forrest E. Linder. Recent trends in health statistics. Science and Public Policy, 3(4):323–331, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Locke:1977:RPS

[Loc77] John H. Locke. Reflections on protecting society. Science and Public Policy, 4(2):160–162, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Locke:1978:SPD

[Loc78] Brian Locke. Science policy and the diffussion of innnovation. Science and Public Policy, 5(4):266– 273, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Lomax:1977:GED

[Lom77] David F. Lomax. Government economic decision making. Science and Public Policy, 4(3):259– 267, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mabbutt:1978:CEA

[Mab78] J. A. Mabbutt. Climatic and ecological aspects of desertification. Science and Public Policy, 5(2): 138–147, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mackay:1975:BRB

[Mac75] A. L. Mackay. Book review: Science and technology in India by A. Rahman, R. N. Bhargava, M. A. Qureshi, Sudarshan Pruthi. Indian Council for Cultural Relations, New Delhi, 1973. pp 236, Rs. 30. Science and Public Policy, 2(3):138–139, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mackay:1978:TSR

[Mac78a] A. L. Mackay. ‘things are in the saddle and ride mankind’. Science and Public Policy, 5(5):398– 399, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mackay:1978:FB

[Mac78b] Alan Mackay. Flying blind. Science and Public Policy, 5(4):315, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mackay:1979:IAS

[Mac79] Alan L. Mackay. Intelligence as an applied science. Science and Public Policy, 6(1):5– 12, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maddock:1975:STI

[Mad75] Sir Ieuan Maddock. Science, technology and industry in the UK. Science and Public Policy, 2(5): 194–211, May 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Maddox:1978:IP

[Mad78] John Maddox. The idea of progress. Science and Public Policy, 5(1):27–39, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mahler:1975:SPN

[Mah75] Halfdan Mahler. Smallpox: The point of no return. Science and Public Policy, 2(4):178, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mahler:1976:HAY

[Mah76] H. Mahler. Health for all by the year 2000. Science and Public Policy, 3(2):122–128, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Malik:1979:TST

[Mal79] F. Bary Malik. Transfer of science and technology. Science and Public Policy, 6(4):289–294, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marwick:1974:WES

[Mar74] M. G. Marwick. Witchcraft and the epistemology of science. Science and Public Policy, 1(11):335– 341, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Marinker:1976:HCB

[Mar76] Marshal Marinker. Health care and basic rights. Science and Public Policy, 3(1):32–36, January 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mayne:1975:SPW

[May75] Alan Mayne. Science policies for a world of crisis. Science and Public Policy, 2(1):37–41, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mazrui:1978:EWE

[Maz78] Ali A. Mazrui. Effects of Western energy conservation. Science and Public Policy, 5(6):402–406, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McNamara:1978:PP

[McN78] Robert S. McNamara. Poverty and population. Science and Public Policy, 5(6):424–435, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McQueen:1977:SPP

[McQ77a] H. J. McQueen. Science policy for the people. Science and Public Policy, 4(4):361–363, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McQueen:1977:GST

[McQ77b] Hugh J. McQueen. Getting in step with technology. Science and Public Policy, 4(2):201– 203, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

McRobie:1979:AT

[McR79] George McRobie. Appropriate technology. Science and Public Policy, 6(4):246–256, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Meyer:1978:ESS

[Mey78] Niels I. Meyer. Energy in a sustainable society: How much electricity do we really need? Science and Public Policy, 5(5):358– 367, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Michalski:1978:PVS

[Mic78] Wolfgang Michalski. Physical versus social scarcity of minerals. Science and Public Policy, 5(6):416– 423, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mikhailov:1979:CST

[Mik79] A. I. Mikhailov. Communicating scientific and technical information. Science and Public Policy, 6(4):256–266, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moravcsik:1978:ASR

[Mor78a] Michael J. Moravcsik. Applied scientific research and the developing countries. Science and Public Policy, 5(2):82–92, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morehouse:1978:DF

[Mor78b] Ward Morehouse. I. doubts and fears. Science and Public Policy, 5(1):2–4, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Morehouse:1979:CHA

[Mor79] Ward Morehouse. Converting hot air into forward motion: NGOs in UNCSTD. Science and Public Policy, 6(4):218–224, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mott:1975:DG

[Mot75] Sir Nevill Mott. Delusions of grandeur. Science and Public Policy, 2(4):179–181, April 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mott:1977:EGA

[Mot77] Sir Nevill Mott. Europe’s giant accelerator. Science and Public Policy, 4(6):571–572, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Moumouni:1974:SEU

[Mou74] Dr Abdou Moumouni. Solar energy utilisation for developing countries. Science and Public Policy, 1(8):207–209, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mesarovic:1974:FOT

[MPG74] Mihajlo Mesarovic, Eduard Pestel, and Maurice Guernier. Four options for tomorrow. Science and Public Policy, 1(9):245–250, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Muhr:1976:EPP

[Muh76] Gerd Muhr. Employment policy: past and future. Science and Public Policy, 3(4):370– 372, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mumford:1974:EEL

[Mum74] Lewis Mumford. Enough energy for life. Science and Public Policy, 1(11):325–330, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Mumfor:1975:NTM

[Mum75] Lewis Mumfor. Next transformation of man. Science and Public Policy, 2(1):11–13, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nashashibi:1978:AFI

[Nas78] Hikmat Sh. Nashashibi. Arab financial institutions and centres: more involvement in regional growth. Science and Public Policy, 5(1):53–56, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ndongko:1978:RML

[Ndo78] Wilfred A. Ndongko. Research as the missing link. Science and Public Policy, 5(6):435–437, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nicholson:1974:NPE

[Nic74] R. L. R. Nicholson. Nuclear power and energy demands. Science and Public Policy, 1(8):201– 207, August 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nelkin:1979:CCR

[NP79] Dorothy Nelkin and Michael Pollak. Consensus and conflict resolution: the polities of assessing risk. Science and Public Policy, 6(5): 307–318, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Nsekela:1977:CPN

[Nse77] Amon J. Nsekela. Cultural policy for national development. Science and Public Policy, 4(1):68– 80, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

ofBessborough:1977:FNS

[oB77] The Earl of Bessborough. French nuclear strategy. Science and Public Policy, 4(1):88–92, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

ofBessborough:1979:LW

[oB79] Earl of Bessborough. A lost world? Science and Public Policy, 6(1):62–69, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Obuder:1979:CRC

[Obu79] Daniel Obuder. Costa Rica — country without an army. Science and Public Policy, 6(3): 195, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

ofCanada:1977:PAL

[oC77] Science Council of Canada. People and agricultural land. Science and Public Policy, 4(6):548–550, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Odhiambo:1979:RCP

[Odh79] Thomas R. Odhiambo. Recommendations for a continuing programme and institutional arrangements. Science and Public Policy, 6(4):216–218, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ogilvy:1976:BNT

[Ogi76] Audrey A. Ogilvy. Bracknell new town: 21 years of growth. Science and Public Policy, 3(2):134– 139, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ojala:1975:IAH

[Oja75] E. M. Ojala. The importance of agriculture to humanity. Science and Public Policy, 2(5):218– 233, May 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Orr:1977:SRE

[Orr77] David Orr. Social risks and the energy option. Science and Public Policy, 4(2):179–183, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pannenborg:1977:IOS

[Pan77] A. E. Pannenborg. Implications for other sectors of industry. Science and Public Policy, 4 (2):162, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Penner:1978:CME

[Pen78] Peter S. Penner. Cynics, martyrs, and energy conservation. Science and Public Policy, 5(2):105– 110, April 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Perry:1977:AFC

[Per77] Walter Perry. Academic freedom and course creation. Science and Public Policy, 4(5):483– 486, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Petana:1977:ECC

[Pet77a] Walter B. Petana. Education in the control of Chagas’ disease. Science and Public Policy, 4(2):191– 195, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Petrovsky:1977:SHS

[Pet77b] Boris Petrovsky. Soviet health service. Science and Public Policy, 4(4):358–361, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Pines:1978:LNP

[Pin78] James M. Pines. Language of nutrition planning. Science and Public Policy, 5(6):457–460, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Preece:1977:SPT

[Pre77] P. F. W. Preece. Sex and the physics teacher. Science and Public Policy, 4(3):281–283, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Price:1974:TEC

[Pri74] Terence Price. A technological economy in a competitive world. Science and Public Policy, 1(11): 344–353, November 1974. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Prigogine:1977:MS

[Pri77] I. Prigogine. Metamorphosis of science. Science and Public Policy, 4(5):397–414, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Prokofiev:1976:DEU

[Pro76] Mikhail Prokofiev. Development of education in USSR. Science and Public Policy, 3(3):261– 268, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1975:TT

[Rah75] Dr A. Rahman. The threat of technology. Science and Public Policy, 2(12):551–552, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rahman:1979:STD

[Rah79] A. Rahman. Science, technology, and development in a new social order. Science and Public Policy, 6(1):12–24, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramphal:1977:HRD

[Ram77a] Shridath S. Ramphal. Human rights and development. Science and Public Policy, 4(3):288– 290, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramphal:1977:NSI

[Ram77b] Shridath S. Ramphal. North/ south issues. Science and Public Policy, 4(4):366–392, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramesh:1979:NFI

[Ram79a] Jairam Ramesh. National financial institutions and technological development. Science and Public Policy, 6(6):405–419, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramphal:1979:CPA

[Ram79b] Shridath S. Ramphal. Commonwealth press as an agent for change. Science and Public Policy, 6(3):188–194, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ramphal:1979:QVH

[Ram79c] Shridath S. Ramphal. ‘quo vadis, homo sentiens?’. Science and Public Policy, 6(2):74–82, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ravetz:1977:TPR

[Rav77] J. R. Ravetz. Turning points in the relations of science with society. Science and Public Policy, 4(5):445–453, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ravetz:1978:ESP

[Rav78] Jerome R. Ravetz. Exploring social phenomena. Science and Public Policy, 5(4):317–318, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ravetz:1979:PPA

[Rav79] J. R. Ravetz. Public perceptions of acceptable risks. Science and Public Policy, 6(5):298– 306, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reddy:1977:TSR

[Red77] A. K. N. Reddy. Technology and self-reliance. Science and Public Policy, 4(3):231–234, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Reeves:1977:CYU

[Ree77] Nancy Reeves. Challenge of the year 2000: 2 ultimate selection. Science and Public Policy, 4(6): 556–557, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rendel:1979:APV

[Ren79] J. Rendel. Animal production in the veterinary curriculum. Science and Public Policy, 6(1):37– 43, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Robertson:1977:PN

[RG77] Jane Robertson and Philip Gummett. Politics of noise. Science and Public Policy, 4(1):59– 63, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Richan:1976:AAP

[Ric76] F. Keith Richan. Adjustment assistance problems in developed countries. Science and Public Policy, 3(3):282–285, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ritterbush:1974:V

[Rit74] Philip C. Ritterbush. viewpoint.

Science and Public Policy, 1(10): 268, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roberts:1975:CCE

[Rob75a] Walter Orr Roberts. Climate change and its effect on world food. Science and Public Policy, 2(6):264–266, June 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roberts:1975:CCA

[Rob75b] Walter Orr Roberts. Climatic change and its affect on world food. Science and Public Policy, 2 (1):26–28, January 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Roberts:1978:FRG

[Rob78] P. C. Roberts. Futures research in government. Science and Public Policy, 5(6):438–443, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rodale:1978:WOC

[Rod78] Robert Rodale. Where are our children’s farms? Science and Public Policy, 5(5):394–397, October 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ronayne:1975:SPP

[Ron75] Jarlath Ronayne. Science Policy planning in Australian Universities: The sixth report of the Universities Commission. Science and Public Policy, 2(9):383–386, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ronayne:1977:AUS

[Ron77] J. Ronayne. Australian universities in a steady state. Science and Public Policy, 4(6):565–569, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rothman:1979:GD

[Rot79a] Harry Rothman. Green dreams. Science and Public Policy, 6(5): 377–378, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Rothman:1979:TSC

[Rot79b] Harry Rothman. Technological small change. Science and Public Policy, 6(5):375–377, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Radice:1977:IS

[RS77] Hugo Radice and Keith Smith. Isolation or subordination? Science and Public Policy, 4(2):203– 206, April 1977. CODEN ????

ISSN 0302-3427 (print), 1471-5430 (electronic).

Sabet:1979:DS

[Sab79] Adel A. Sabet. Democratization of science. Science and Public Policy, 6(5):371–375, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sagasti:1975:SRN

[Sag75a] F. R. Sagasti. Self-reliance 1: Notes on technological selfreliance and cooperation among Third World countries. Science and Public Policy, 2(10):416–424, October 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sagasti:1975:NPS

[Sag75b] Francisco R. Sagasti. Notes on the progress of the STPI project. Science and Public Policy, 2(9):380– 382, September 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sagasti:1977:GTP

[Sag77a] Francisco R. Sagasti. Guidelines for technology policies. Science and Public Policy, 4(1):2– 15, February 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sagasti:1977:RST

[Sag77b] Francisco R. Sagasti. Reflections on scientific-technological revolutions. Science and Public Policy, 4(6):504–513, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sagasti:1979:STP

[Sag79] Francisco R. Sagasti. The science and technology policy instruments project. Science and Public Policy, 6(4):275–288, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salomon:1977:CSS

[Sal77a] Jean-Jacques Salomon. Contribution of the social sciences. Science and Public Policy, 4(4):307– 318, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Salomon:1977:CSC

[Sal77b] Jean-Jacques Salomon. Crisis of science, crisis of society. Science and Public Policy, 4(5):414– 433, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Seamans:1975:WTT

[Sea75] Robert C. Seamans, Jr. Western technology for the Third World. Science and Public Policy, 2(3): 137, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Semenov:1977:HHW

[Sem77] Anatoly Semenov. Health hazards in work environment. Science and Public Policy, 4(3):242– 243, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sherwell:1977:WSP

[She77] Chris Sherwell. Wanted: a science policy. Science and Public Policy, 4(4):339–341, August 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Siebker:1975:CRP

[Sie75] Manfred Siebker. Club of rome: a progress report. Science and Public Policy, 2(3):122–123, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Shang-Kuei:1975:WCC

[SK75] Chen Shang-Kuei. Water conservation in China. Science and Public Policy, 2(12):548–550, December 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Soen:1978:PAI

[Soe78] Dan Soen. Purposes and achievements of Israeli agricultural research. Science and Public Policy, 5(6):460–469, December 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Soyinka:1978:TA

[Soy78] Wole Soyinka. Technology and the artist. Science and Public Policy, 5(1):65–66, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Stafford:1977:SBV

[Sta77] J. Stafford. Safety balance: VCM case study. Science and Public Policy, 4(2):134–142, April 1977.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Sudarskis:1979:NCS

[Sud79] Michel Sudarskis. NATO’s contribution to science, technology, and development. Science and Public Policy, 6(4):241–246, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Szczepanik:1977:RCW

[Szc77] Edward F. Szczepanik. Recent changes in world livestock/feed price ratios. Science and Public Policy, 4(3):255–257, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Talbot:1979:FTF

[Tal79] James J. Talbot. The fate of tropical forests. Science and Public Policy, 6(3):185–188, June 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teitel:1978:CAT

[Tei78] Simón Teitel. Concern for appropriate technology. Science and Public Policy, 5(1):64–65, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teller:1977:PPN

[Tel77] Edward Teller. Pollution by poverty: the need for nuclear energy. Science and Public Policy, 4 (3):234–238, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thapalyal:1976:MLA

[Tha76] Lalit Thapalyal. Mongolia: a leap across the centuries. Science and Public Policy, 3(4):331– 338, August 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Thorsson:1974:CTG

[Tho74] Inga Thorsson. Communication technology in a global context. Science and Public Policy, 1(12): 396–397, December 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tinbergen:1976:NIO

[Tin76] Jan Tinbergen. A new international order. Science and Public Policy, 3(2):128–133, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Tittmar:1979:MMN

[Tit79] H-G Tittmar. Multistat man: the next step in computer evolution. Science and Public Policy, 6(2): 85–90, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Toren:1978:IPI

[Tor78] Nina Toren. Israeli policy for integrating Soviet scientists. Science and Public Policy, 5(3):181– 190, June 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Toren:1979:SWS

[Tor79] Nina Toren. Scientific work in the Soviet union: insights from emigré scientists in the USA. Science and Public Policy, 6(6):386–395, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Townsend:1974:BBB

[Tow74] H. Townsend. Big business and big science. Science and Public Policy, 1(10):290–297, October 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Trip:1976:ICS

[Tri76] Fokele P. H. Trip. International cooperation and social sciences in a technological Europe. Science and Public Policy, 3(2):139– 142, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Unesco:1979:STM

[Une79] Paris Unesco. Scientific and technical manpower resources. Science and Public Policy, 6(6):432–443, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Ustinov:1977:CYI

[Ust77]  Peter Ustinov.  Challenge of

the year 2000: 1 the individual and the collective. Science and Public Policy, 4(6):552–556, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Vaizey:1977:SEI

[Vai77] Lord Vaizey. Socio-economic implications. Science and Public Policy, 4(2):162–163, April 1977.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanDam:1979:ESC

[vD79] André van Dam. Energy saved is cheaper than energy generated. Science and Public Policy, 6(1):2– 5, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Verghese:1977:EDJ

[Ver77] B. George Verghese. Experiment in development journalism. Science and Public Policy, 4(6):513– 517, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vonHippel:1979:AIA

[vHF79] Eric von Hippel and Stan N.

Finkelstein. Analysis of innovation in automated clinical chemistry analyzers. Science and Public Policy, 6(1):24–37, February 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vonMoltke:1977:AEC

[vM77] Heinrich von Moltke. Attitude of the European commission. Science and Public Policy, 4(2):148– 159, April 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

vanTeutem:1976:GSP

[vT76] Onno van Teutem. Grain stock policies in developing countries. Science and Public Policy, 3(3): 207–214, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weinberg:1976:URI

[Wei76] Dr Alvin M. Weinberg. The university, the research institutions and society. Science and Public Policy, 3(2):103–107, April 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weingart:1977:STE

[Wei77] Peter Weingart. Science today and its environment. Science and Public Policy, 4(5):??, October 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Weiss:1979:MTD

[Wei79] Charles Weiss, Jr. Mobilizing technology for developing countries. Science and Public Policy, 6(5):359–371, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wheeler:1979:EOS

[Whe79] John Archibald Wheeler. Einstein and other seekers of the larger view. Science and Public Policy, 6(6):396–404, December 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams:1976:MTS

[Wil76] Maurice J. Williams. Mid-term in the second development decade. Science and Public Policy, 3(3): 252–261, June 1976. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilkes:1977:TL

[Wil77] A. C. Wilkes. Technology as a language. Science and Public Policy, 4(6):??, December 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilkes:1978:TR

[Wil78a] A. C. Wilkes. Technological rights.

Science and Public Policy, 5(4): 274–279, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilson:1978:LGC

[Wil78b] Richard Wilson. “Let’s get on with cutting the cake… ”. Science and Public Policy, 5(1):40– 41, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wilkes:1979:WKB

[Wil79a] A. C. Wilkes. Whizz kids and boffins. Science and Public Policy, 6(5):338–345, October 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Williams:1979:NWP

[Wil79b] I. G. K. Williams. Nuclear wastes:

a problem of perspective? Science and Public Policy, 6(2):134– 142, April 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Winthrop:1974:IPP

[Win74] Henry Winthrop. Interdisciplinary policy planning. Science and Public Policy, 1(11):353–359, November 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Walton:1979:WST

[WL79] Barbara Walton and Mildred Robbins Leet. Women in science and technology for development. Science and Public Policy, 6(4):225– 229, August 1979. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Wrigley:1978:EEP

[Wri78] E. A. Wrigley. Estimating England’s population before the first census. Science and Public Policy, 5(4):290–295, August 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Yaacov:1975:AIS

[Yaa75] Professor I. Yaacov. Arab–Israeli science and technology gap. Science and Public Policy, 2(3):135– 136, March 1975. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zabrodin:1978:RBM

[Zab78] Yuri Zabrodin. Relationship between man and the machine. Science and Public Policy, 5(1):44– 46, February 1978. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Zahlan:1974:AIS

[Zah74] A. B. Zahlan. Arab–Israeli science and technology gap. Science and Public Policy, 1(9):239–245, September 1974. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).